### v1

# FedLog: Personalized Federated Classification with Less Communication and More Flexibility

Abstract
In federated learning (FL), the common paradigm that FedAvg proposes and most algorithms follow is that clients train local models with their private data, and the model parameters are shared for central aggregation, mostly averaging. In this paradigm, the communication cost is often a challenge, as modern massive neural networks can contain millions to billions parameters. We suggest that clients do not share model parameters but local data summaries, to decrease the cost of sharing. We develop a new algorithm FedLog with Bayesian inference, which shares only sufficient statistics of local data. FedLog transmits messages as small as the last layer of the original model. We conducted comprehensive experiments to show we outperform other FL algorithms that aim at decreasing the communication cost. To provide formal privacy guarantees, we further extend FedLog with differential privacy and show the trade-off between privacy budget and accuracy.
1 Introduction
This work focuses on the design of techniques that are communication efficient and flexible for personalized federated learning (FL). The main challenge in FL is to enable communication between different clients and a server to facilitate the estimation of (personalized) aggregated models while keeping private the data in each client [1]. Hence, instead of sharing the data into a central location, clients share parameters of their local model which are then aggregated. However, this may induce large communication costs when the models are large (millions to billions of parameters) [1] and most aggregation schemes require homogeneous architectures across clients [2, 3, 4, 5, 6]. While the data never leaves the clients, the shared parameters may be thought as implicit data summaries since they indirectly capture information about the data. Given this perspective, one may ask: can we reduce communication costs by sharing small data summaries while preserving privacy and can clients utilize different model architectures?
This paper describes a new algorithm called FedLog that shares data summaries with much smaller dimensionality than model parameterization. Furthermore, clients may use different model architectures to compute those summaries. Since the parameters of those local models are never shared, they do not need to use the same architecture. Data summaries are aggregated in a central server that trains a global prediction head by Bayesian inference. More specifically, it treats the data summaries as sufficient statistics in an exponential family model and returns to the clients the canonical parameters of this exponential family model. While all clients use this same head for prediction, their locally trained models for data summarization enable them to personalize their predictions. Differential privacy (DP) is incorporated to mitigate privacy leakage via the data summaries. The paper makes the following contributions:
  • FedLog: a new Federated Bayesian logistic regression algorithm, the model of which is carefully designed to provide a statistical interpretation for representation learning.
  • Experiments demonstrating FedLog’s low communication cost (as small as 1.6% of the original model per round) and fast convergence under multiple situations.
  • Incorporation of DP and demonstration of favorable trade-off between privacy and utility.
2 Background
2.1 Federated Learning and Personalized FL
In Federated Learning (FL), there is always a collection of clients $c\in S$ who wish to collaborate, and each client holds their own data $\mathcal{D}_{c}=({\bf X}_{c},{\bf y}_{c})$ locally. We seek to train some machine learning model $\mathcal{M}_{\theta}$ with parameters $\boldsymbol{\theta}$ on these client data. Conventionally, we would centralize all the data $\mathcal{D}=\bigcup_{c\in S}\mathcal{D}_{c}$, and choose some algorithm $\mathcal{A}$ and parameter initialization $\boldsymbol{\theta}_{0}$ to learn the model:
$\mathcal{M}_{\boldsymbol{\theta}}=\mathcal{A}(\mathcal{D},\boldsymbol{\theta}_{0})$
This simple approach becomes infeasible if the clients cannot directly share their data due to privacy concerns. FL intends to tackle this problem. A trusted central server is allowed to receive and send perturbed matrices that only contain limited information about raw client data, such as model parameters. Most FL algorithms follow the following paradigm. Such algorithms iterate between local optimization and global aggregation. During local optimization at $t^{th}$ iteration, all the clients (or a proportion of the clients in some cases) start with the global consensus parameters $\theta^{t}$, and run the algorithm $\mathcal{A}$ to update the local models: $\mathcal{M}_{\boldsymbol{\theta}^{t+1}_{c}}=\mathcal{A}(\mathcal{D}_{c},\boldsymbol{\theta}^{t})$. Note the local models could have other local parameters $\tilde{\boldsymbol{\theta}}$ than $\boldsymbol{\theta}$, but is omitted in the notation for simplicity. Then, clients potentially compress their updated model parameters, and send them instead of data to the central server so that it can be aggregated into the global consensus parameters for next iteration $\boldsymbol{\theta}^{t+1}$.
In this paradigm, one challenge is how to aggregate model parameters so that the resulting global consensus $\boldsymbol{\theta}^{t+1}$ is a better approximation to the centralization version. This is especially challenging when the number of local training epochs $>1$ [7], due to non-linear loss functions and predictors, thus many FL algorithms simply resort to the averaging heuristic [2, 3, 4, 5, 6]. For example, the very first FL algorithm, FedAvg [2], let the clients send all model parameters to the server without compressing, and their weighted average becomes the new global consensus.
Another challenge is that clients often have different data distribution $\Pr({\bf X}),\Pr({\bf y})$ or even $\Pr({\bf y}|{\bf X})$. This is referred to as non-i.i.d. clients, and it could harm the global aggregation in terms of convergence rate and model utility [8], especially when the goal is to learn a global model. Thus, personalized FL is introduced, where a global model is not mandatory, but each client could have their own model that best fit their data distribution. See [9] for more motivation and algorithms.
2.2 Exponential Family
Exponential family refers to a set of probability distributions of the following canonical form.
$\Pr({\bf x}|\boldsymbol{\eta})=h({\bf x})\exp(\boldsymbol{\eta}^{\top}{\bf T}({\bf x})-A(\boldsymbol{\eta}))$ (1)
[…]
2.3 Conjugate Priors
Suppose we wish to model the distribution of some data ${\bf x}$ by learning some parameters ${\boldsymbol{\eta}}$, and denote the likelihood induced by the model $\Pr({\bf x}|\boldsymbol{\eta})$. With Bayesian inference, $\boldsymbol{\eta}$ is often treated as random variables, and a prior distribution $\Pr(\boldsymbol{\eta})$ can be specified to incorporate prior knowledge. Then, by Bayes’ Theorem the posterior distribution is:
[…]
The integral in Eq. 2 and thus the posterior $\Pr(\boldsymbol{\eta}|{\bf x})$ may not be tractable for arbitrary priors $\Pr(\boldsymbol{\eta})$. A convenient choice that guarantees analytical solutions is the conjugate prior. Given a likelihood $\Pr({\bf x}|\boldsymbol{\eta})$, a prior is called its conjugate prior if $\Pr(\boldsymbol{\eta})$ and $\Pr({\bf x}|\boldsymbol{\eta})$ follow the same distribution family. Specifically, if the likelihood is an exponential family, it has known conjugate priors [11].
Likelihood: $\Pr({\bf x}|\boldsymbol{\eta})=h({\bf x})\exp(\boldsymbol{\eta}^{\top}{\bf T}({\bf x})-A(\boldsymbol{\eta}))$ (3)
Prior: $\Pr(\boldsymbol{\eta};\boldsymbol{\chi},\nu)=f(\boldsymbol{\chi},\nu)\exp(\boldsymbol{\eta}^{\top}\boldsymbol{\chi}-{\nu}A(\boldsymbol{\eta}))$ (4)
Posterior: $\Pr(\boldsymbol{\eta}|{\bf x})=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+{\bf T}({\bf x}),\nu+1)$ (5)
[…]
2.4 Related Work
Representation learning.
The most related to our works are federated representation learning methods. These approaches separate a whole neural network into two parts, body and head. They interpret the body as a deep neural network that learns a compact feature representation from raw data, and the head as a few shallow layers that map the features to the prediction. LG-FedAvg [3] chooses to localize the body and learn local feature representations, but averages the head for global information sharing. On the contrary, FedPer [4] localizes the head but averages the body to learn a universal representation. Later, FedRep [5] also localizes the head, but instead of jointly updating the body and the head locally, they first update the local head to convergence, and then fix the head and update the body. FedProto [12] localizes the whole model, but averages feature representations by class and forces local models to learn similar representations. FedLog can also be interpreted as a representation learning algorithm, where we learn local representations by all the layers except the last one. However, unlike previous works that share model parameters and heuristically average them, we share data summaries and update the global head with Bayesian inference.
Another line of works (CCVR [13], FedPFT [14]) fit Gaussian Mixture Models (GMM) to local features extracted by a globally uniformed body, and shares the natural parameters. The server then draws virtual features from the GMM and train a global model. It is notable that these algorithms rely on pretrained FedAvg or foundation models to unify the body. Although FedLog also fits a exponential family distribution to local features, we do not need such pretrained models since our bodies are trained locally with any architecture. Also, we work with the canonical parameters (which result in smaller messages) and do not need to draw virtual samples from the learnt distribution.
Communication efficient FL.
In addition to localizing part of the original model, which decreases the message size straightforwardly, other lines of work that aim at communication efficient FL include optimization algorithms, client selection, and model compression [1]. First, it has been shown that local training epochs could impact the rounds of global communication needed [2]. Based on this observation, researchers proposed different local optimization methods to reduce the communication rounds [15, 16, 17]. On the other hand, instead of updating all clients or randomly selecting a subset of clients, researchers proposed various algorithms that accelerate the global learning process by carefully selecting clients that contribute more to the global model or are more efficient in terms of uploading time [18, 19, 20, 21]. Additionally, researchers proposed algorithms that transmit smaller messages than the whole model by reducing and compressing the parameters [22, 23, 24]. However, most such algorithms are efficient at the cost of model accuracy [25].
Bayesian FL.
Other Bayesian models have been explored to represent distributions over models and predictions in FL. The challenge is in the aggregation of the local posteriors. Various techniques have been proposed including personalized GPs [6], posterior averaging [26], online Laplace approximation [27], Thompson sampling [28], MCMC sampling [29]. There is not much in common between these techniques with our approach despite the use of Bayes theorem since prior Bayesian techniques tend to emphasize on calibration, approximating the posterior, or even different tasks.
3 Federated Bayesian Logistic Regression
In this section, we first detail our model construction and algorithm. Then, we show how to incorporate differential privacy to prevent potential privacy leakage.
3.1 FedLog
Notations.
Let $\tilde{\boldsymbol{\theta}}_{c}$ denote the parameters of a local neural network of client $c$. Let $f_{\tilde{\boldsymbol{\theta}}_{c}}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{m}$ be the function of the local neural network and $\boldsymbol{\Phi}_{c}=f_{\tilde{\boldsymbol{\theta}}_{c}}({\bf X}_{c})$ be the $m$ local features extracted from the local input. For convenience, we designate one of the features to be always $1$. Let $n_{class}$ denote the total number of classes in a classification task, and $n_{c}$ denote the size of the local dataset. Let ${\bf e}_{i}\in\mathbb{R}^{n_{class}}$ be the standard basis (i.e. one-hot vector form) of label $i\in\{1,2,\cdots,n_{class}\}$, and $\otimes:\mathbb{R}^{m}\times\mathbb{R}^{n_{class}}\rightarrow\mathbb{R}^{m*n_{class}}$ be the Kronecker product.
We assume the joint probability of each data point in $(\boldsymbol{\Phi}_{c},{\bf y}_{c})$, denoted as $\Pr(\boldsymbol{\phi},y)$, is an exponential family with canonical parameters $\boldsymbol{\eta}\in\mathbb{R}^{m*n_{class}}$ as follows.
sufficient statistics: ${\bf T}(\boldsymbol{\phi},y):=\boldsymbol{\phi}\otimes{\bf e}_{y}$ (6)
base measure: $h(\boldsymbol{\phi},y):=\frac{\exp\left(-\sum_{i=1}^{m}\boldsymbol{\phi}_{i}^{2}\right)}{\sqrt{\pi^{m}}}$ (7)
where $\boldsymbol{\phi}_{i}\in\mathbb{R}$ denotes the $i^{th}$ entry of $\boldsymbol{\phi}$. One advantage with this specific choice of $\bf T$ and $h$ is that $A$ has an explicit expression. Let $\boldsymbol{\eta}_{y}\in\mathbb{R}^{m}$ denote the $((y-1)*m)^{th}$ to $(y*m)^{th}$ entries of $\boldsymbol{\eta}$, $y\in\{1,2,\cdots,n_{class}\}$. Let $\boldsymbol{\eta}_{y,i}\in\mathbb{R}$ denote the $i^{th}$ entry of $\boldsymbol{\eta}_{y}$. Then,
[…]
We can write out the likelihood, conditional likelihood, prior, and posterior as follows.
$\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})$ $=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}-\boldsymbol{\phi}^{\top}\boldsymbol{\phi})}{\sqrt{\pi^{m}}\sum_{y^{\prime}=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\eta}_{y^{\prime}}/4)}$ (9)
$\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$ $=\frac{\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})}{\Pr(\boldsymbol{\phi}|\boldsymbol{\eta})}=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi})}{\sum_{y^{\prime}=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\phi})}$ (10)
$\Pr(\boldsymbol{\eta};\boldsymbol{\chi},\nu)$ $=f(\boldsymbol{\chi},\nu)\frac{\exp(\boldsymbol{\eta}^{\top}\boldsymbol{\chi})}{(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\eta}_{y^{\prime}}/4))^{\nu}}$ (11)
$\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},y)$ $=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\phi}\otimes{\bf e}_{y},\nu+1)$ (12)
Unfortunately, the normalizing constant function $f(\boldsymbol{\chi},\nu)$, the marginal likelihood $\Pr(y|{\boldsymbol{\phi}})$, and the predictive posterior $\Pr(y_{*}|\boldsymbol{\phi}_{*},\boldsymbol{\phi},y)$ are still intractable. However, another advantage of our choice of ${\bf T}$ and $h$ is shown by Eq. 10: the conditional likelihood is exactly the softmax function over $\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}$. This means we can take any deep neural network that extracts $m$ features, and append $\boldsymbol{\eta}$ as the last linear layer that maps the features to $n_{class}$ logits. Then, this composed neural network serves as a stand-alone classifier that computes the conditional likelihood $\Pr(y|{\boldsymbol{\phi}},\boldsymbol{\eta})$.
Algorithm 1 FedLog (${\bf X}_{c},{\bf y}_{c}$: local data, $\tilde{\boldsymbol{\theta}}_{c}$ local body parameters, $\boldsymbol{\eta}$: global head parameters, $\boldsymbol{\chi}$: prior parameter, $\nu$: prior parameter, $\zeta$: local learning rate)
  Server: initializes $\boldsymbol{\eta}$
  for each client $c\in S$ do
     Initialize $\tilde{\boldsymbol{\theta}}_{c}$
  for each global update round do
     Server: sends $\boldsymbol{\eta}$ to clients
     for each client $c\in S$ do
        for each local update round do
           $\tilde{\boldsymbol{\theta}}_{c}\leftarrow\tilde{\boldsymbol{\theta}}_{c}-\zeta\nabla\mathcal{L}_{c}$ (Eq. 13)
        $\boldsymbol{\Phi}_{c}\leftarrow f_{\tilde{\boldsymbol{\theta}}_{c}}({\bf X}_{c})$
        Send $\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},n_{c}$ to server
     Server:
     $\boldsymbol{\Phi}\leftarrow\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},n\leftarrow\sum_{c\in S}n_{c}$
     $\boldsymbol{\eta}\leftarrow\operatorname*{arg\,max}_{\boldsymbol{\eta}}\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)$ (Eq.  15)
Based on the current model, we propose a new algorithm named FedLog, as shown in Alg. 1. At the beginning, the server initializes $\boldsymbol{\eta}$ (the global head) randomly, and the clients initializes $\tilde{\boldsymbol{\theta}}_{c}$ (the local bodies) either completely randomly, or with the same random seed sent by the server to unify the initialization. Note $\tilde{\boldsymbol{\theta}}_{c}$ is not part of our exponential family assumption, thus we do not require them to have the same shape or architecture amongst different clients.
Parameters that we need to optimize are essentially $\tilde{\boldsymbol{\theta}}_{c}$ and $\boldsymbol{\eta}$, which can be done by maximizing $\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$ and $\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},{y})$ in turns iteratively, similarly to the expectation-maximization algorithm. Concretely, all the clients $c\in S$ first fix the global head $\boldsymbol{\eta}$, and update their local bodies $\tilde{\boldsymbol{\theta}}_{c}$ with gradient descent. We derive the loss function as:
$\begin{split}\mathcal{L}_{c}&=-\sum_{i=1}^{n_{c}}\ln\Pr({\bf y}_{c,i}|\boldsymbol{\Phi}_{c,i},\boldsymbol{\eta})=-\sum_{i=1}^{n_{c}}\ln\frac{\exp(\boldsymbol{\eta}_{y_{c,i}}^{\top}\boldsymbol{\Phi}_{c,i})}{\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\Phi}_{c,i})}\end{split}$ (13)
where $(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})\in\mathbb{R}^{m}\times\mathbb{R}$ is the $i^{th}$ data point of $\boldsymbol{\Phi}_{c},{\bf y}_{c}$. This is exactly the cross entropy loss widely used in deep learning for classification tasks. Then, clients compute sufficient statistics of their local data $\sum_{i=1}^{n_{c}}{\bf T}(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})=\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}}$. They then send the summations and $n_{c}$ to the server for global head learning. As we have discussed in Sec. 2.2, the sufficient statistics contain all the information in the raw dataset that could be used to infer $\boldsymbol{\eta}$ in our model. In this case, the server only needs to know the summation of all the sufficient statistics, since
$\begin{split}&\Pr(\boldsymbol{\eta}|\boldsymbol{\Phi}_{c_{1}},{\bf y}_{c_{1}},\cdots,\boldsymbol{\Phi}_{c_{k}},{\bf y}_{c_{k}})=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},\nu+\sum_{c\in S}n_{c})\end{split}$ (14)
can be trivially inferred from Eq. 12. Note the size of the message sent by each client equals the size of the last linear layer. After receiving the sufficient statistics, the server computes $\boldsymbol{\Phi}=\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},n=\sum_{c\in S}n_{c}$ and updates the global head $\boldsymbol{\eta}$ by maximum a posteriori (MAP):
$\begin{split}\boldsymbol{\eta}&=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\frac{\exp\left(\boldsymbol{\eta}^{\top}(\boldsymbol{\chi}+\boldsymbol{\Phi})\right)}{(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\eta}_{y}/4))^{(\nu+n)}}\end{split}$ (15)
Note Eq. 15 is a convex optimization task, and we can easily compute its sole maximum by gradient descent. Then, the server sends $\boldsymbol{\eta}$ back to all the clients and starts the next round of updates. The process is repeated until convergence.
3.2 Interpret FedLog
In this section, we analyze our assumptions in more details and give some insights about how FedLog works. The first assumption we made is that $\forall c\in S$, the local data points $\boldsymbol{\Phi}_{c},{\bf y}_{c}$ are from the same exponential family distribution whose pdf is given by Eq. 9. In other words, we assume that the local bodies $\tilde{\boldsymbol{\theta}}_{c}$ transform their input, of any form, to the same representation space. This may seem infeasible at the first glance, since we do not directly aggregate the local body parameters, and they may take any architecture. However, note we fix the global head $\boldsymbol{\eta}$ during local updates, by which the local bodies are forced to learn a universal representation space. See Sec. 4.1 for more details. This assumption allows principled discriminative training for the local bodies with cross entropy loss, but unavoidably leaves a generative model for learning the global head. We can further see that $\Pr(\boldsymbol{\phi}|y,\boldsymbol{\eta})\propto\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}-\boldsymbol{\phi}^{\top}\boldsymbol{\phi})$, which is the kernel of a multivariate Gaussian distribution. This means we essentially assumed a mixture of Gaussians for the local features $\boldsymbol{\phi}$. This is a mild assumption, since sufficiently expressive DNN encoders can map data to any distribution any closely.
The second assumption we made is the prior of the global head. Since $\boldsymbol{\eta}$ has a support over $\mathbb{R}^{m*n_{class}}$, it is impossible to specify a uniform prior. Without any prior knowledge, we can set ${\boldsymbol{\chi}}={\bf 0},\nu=1$. The prior then becomes $\Pr(\boldsymbol{\eta})\propto\exp(-A(\boldsymbol{\eta}))=(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\eta}_{y}/4))^{-1}$. The pdf takes its maximum at $\boldsymbol{\eta}={\bf 0}$ and decreases quickly as the absolute value of entries of $\boldsymbol{\eta}_{y}$ grows larger. This is in analogy to the Lasso regularizer in the regression case, which prevents the model from learning coefficients with large absolute values that are more likely due to noise or over-fitting.
From the Bayesian view, FedLog can take any deep classifier, and make the last linear layer Bayesian. It essentially operates a Bayesian logistic regression model on the local representations, thus the name FedLog. We start from a generative assumption and achieve the cross entropy loss conditional likelihood usually assumed directly in Bayesian logistic regression. We obtained an analytical solution for the kernel of the posterior, which can be calculated easily by the summation of sufficient statistics, whose sizes are as small as the size of the last linear layer, and avoid the need to share separated data points.
From the federated representation learning view, FedLog has a one-layer global head and a deep local body. It iterates between learning local representations and learning global linear separators as if it has seen all the local representations. The two learning processes are completely separated, unlike the common paradigm where the local representations and the linear separators are often optimized jointly. The clients are only responsible for moving local representations to the correct sides of the fixed linear separator. The server is only responsible for finding the best linear separator given the local representations. As we will show with the experiments, we can automatically learn universal local representations and converge faster than using the averaging heuristic.
3.3 Differential Privacy
FL can be combined with formal mechanisms such as differential privacy (DP) [30, 31] or secure multi-party computation (MPC) [32, 33, 34], to provide formal privacy guarantees. We now extend FedLog to be differentially private.
Theorem 3.1.
If the absolute value of features are clipped to $b$, FedLog messages satisfy $(\epsilon,\delta)$-DP with additive Gaussian noise ${\bf T}^{\prime}(\boldsymbol{\Phi}_{c},{\bf y}_{c}):={\bf T}(\boldsymbol{\Phi}_{c},{\bf y}_{c})+\mathcal{N}({\bf 0},\sigma^{2}{\bf I})$, where $\sigma=\sqrt{2(1+(m-1)*b^{2})\ln(1.25/\delta)}/{\epsilon}$.
Proof.
$L_{2}({\bf T})=\max_{\mathcal{D},\mathcal{D}^{\prime}}||{\bf T}(\mathcal{D})-{\bf T}(\mathcal{D}^{\prime})||_{2}=\max_{\boldsymbol{\phi},y}||{\bf T}(\boldsymbol{\phi},y)||_{2}=\sqrt{1+(m-1)*b^{2}}$
See more details in Appendix A. ∎
When $n$ is large enough, the amount of noise (independent of $n$) becomes negligible to the model. Optionally, with secure MPC, it is sufficient to add such Gaussian noise once globally to the summation of all the messages each global update round. The server then computes the posterior normally with the noisy sufficient statistics.
4 Experiments
4.1 Synthetic
We designed the following synthetic experiment to justify our claim that FedLog can learn universal local representation spaces without sharing local feature extractors, and argue why FedLog can converge faster than prior arts. As shown in the top left image of Fig. 1, we first sample $80$ two-dimensional training data points $(x_{1},x_{2})$ uniformly from the $[-5,5]\times[-5,5]$ square. Data points are separated into class 0 (blue dots) and class 1 (green dots) by the circle located at the origin and of radius $26/7$. These data points are further divided evenly into two sets, client 0 (dark dots) and client 1 (light dots), based on ordered $x_{1}$ to simulate non-i.i.d. clients. Client 0 operates a three-layer fully connected feature extractor, while client 1 operates a two-layer fully connected feature extractor, to simulate clients with different computational resources and showcase the flexibility. We compare FedLog to LG-FedAvg with a one-layer global head, in which case the size of shared messages is the same between those two algorithms. The top middle image of Fig. 1 shows the local representations and linear separators with random initialization. Dashed lines are the linear separators induced by the global head. To be fair, the head of both clients are initialized to be the same for LG-FedAvg. We run FedLog and LG-FedAvg for one global update round, with $1$ to $30$ local iterations. The bottom left and middle images of Fig. 1 respectively show the models LG-FedAvg and FedLog converge to locally. LG-FedAvg learns very different local representations and linear separators even if the last layer is initialized to be the same, and the averaged global linear separator (red dashed line) is clearly suboptimal. This is because it jointly updates the feature extractor and the linear separator, and the local updates diverge in different directions. On the contrary, FedLog clients learn universal local representations with fixed last linear layer, and the server is able to draw the linear separator as if it has seen all client data. Finally, the top and bottom right images show the training and testing accuracy v.s. local iterations. The testing data points are sampled i.i.d. from the same distribution. The results show: i) FedLog makes more progress in one global update round than other averaging based prior arts; ii) FedLog is resistant to over-fitting, as the difference between training and testing accuracy is small; iii) FedLog can learn universal local representations by fixing the last layer, even with different initialization and architectures of the feature extractor.
Figure 1: Synthetic experiments. Dots are data points or local representations: dark green: client 0 class 0; light green: client 1 class 0; Dark blue: client 0 class 1; light blue: client 1 class 1. Dashed lines are linear separators. Accuracy results are averaged over 6 seeds.
Table 1: Testing accuracy and communication cost reported for MNIST, CIFAR10, and CIFAR100. Accuracy reports the mean $\pm$ standard error of the testing accuracy over $10$ seeds. Higher is better. Communication cost reports the number of bits transmitted between one client and the server each global update round. Lower is better. $\Uparrow$ denotes significantly better results with $p<0.01$; $\Downarrow$ denotes significantly worse results with $p<0.01$.
MNIST CIFAR10 CIFAR100
accuracy comm cost accuracy comm cost accuracy comm cost
FedAvg 89.76$\pm$0.69 698880 26.29$\pm$0.44 4102528 13.34$\pm$0.15 4684288
LG-FedAvg 1 97.85$\pm$0.05 16320 86.57$\pm$0.29 64640 55.00$\pm$0.26 646400
LG-FedAvg 2 98.18$\pm$0.06 529920 85.56$\pm$0.32 839040 54.90$\pm$0.24 1420800
FedPer 96.16$\pm$0.19 168960 83.54$\pm$0.40 3263488 52.82$\pm$0.21 3263488
FedRep 95.51$\pm$0.29 168960 82.96$\pm$0.35 3263488 48.70$\pm$0.29 3263488
CS-FL 79.65$\pm$1.22 72088 23.60$\pm$1.08 423092 4.52$\pm$0.15 966172
FedBabu 86.30$\pm$1.04 682560 25.37$\pm$0.44 4037888 9.70$\pm$0.16 4037888
FedProto 98.19$\pm$0.06 16320 87.37$\pm$0.26 64640 55.32$\pm$0.19 646400
FedDBE 96.79$\pm$0.34 700480 72.77$\pm$0.79 4108928 36.67$\pm$0.85 4690688
FedLog (ours) 98.15$\pm$0.05 16320 87.08$\pm$0.22 64640 56.46$\pm$0.27 646400
Table 2: Lower is better. $\Uparrow$ denotes significantly worse results with $p<0.01$.
MNIST CIFAR10 CIFAR100
rounds needed rounds needed rounds needed
LG-FedAvg 1 5.9$\pm$0.38 40.3$\pm$3.09 67.1$\pm$2.76
FedProto 3.7$\pm$0.21 28.5$\pm$2.05 46.6$\pm$1.90
FedLog(ours) 3.9$\pm$0.38 21.5$\pm$2.23 36.8$\pm$1.39
4.2 Communication Cost
We conduct experiments on MNIST, CIFAR10, and CIFAR100, to show FedLog achieves better accuracy with less communication. We compare FedLog with the following baselines: i) FedAvg [2], which averages the whole model each global update round; ii) LG-FedAvg [3], which localizes bodies and averages heads. Two variants are reported: LG-FedAvg 1 which maintains one global layer and LG-FedAvg 2 which maintains two global layers; iii) FedPer [4], which localizes heads and averages bodies; iv) FedRep [5], which also localizes heads and averages bodies, but trains local heads and bodies separately; v) CS-FL [24], a model compression technique that compresses messages with the compressed sensing framework; vi) FedBabu [35], which averages bodies and never updates heads; vii) FedProto [12], which averages local feature representations by class and forces local models to learn similar representations; viii) FedDBE [36], which averages both bodies and heads but accelerates convergence by learning a domain representation bias. Convolutional neural networks of the same architecture and initialization is used for all the algorithms. FedPer and FedRep localizes the last two fully connected layers for less communication cost.
We first distribute the training set evenly into $50$ (for MNIST) or $100$ (for CIFAR10 and CIFAR100) clients. Each client takes only $2$ classes to maximize the heterogeneity. The testing set is distributed similarly to the clients, following the same distribution as the training data. Specially, for MNIST only, to simulate the difficult situation where clients do not have sufficient local data, we train on only $5\%$ of the training set, while test on the whole testing set. All the clients start from the same initialization, and all the algorithms are run for $100$ (for MNIST and CIFAR10) or $150$ (for CIFAR100) global update rounds. The testing accuracy is recorded after each round. Hyperparameters are optimized beforehand through grid search. See Appendix B for the model architectures, hyperparameters used by each algorithm and other experiment details.
We report the mean and the standard error of the testing accuracy resulting from $10$ seeds in Table 1. We measured the statistical significance of the results compared to FedLog with one-tailed Wilcoxon signed-rank tests [37]. We also report the communication cost of each global update round, namely the number of bits transmitted between one client and the server. The results show FedLog can achieve statistically significant accuracy improvement or at least comparable utility compared to prior arts, with the least communication cost as small as $2.3\%$, $1.6\%$, and $13.8\%$ of the communication cost of FedAvg for MNIST, CIFAR10, and CIFAR100 respectively.
We further report how many global rounds are needed to reach a reasonable accuracy threshold in Table 2. The thresholds are 97%, 83%, and 53% respectively. These numbers are the largest integer of accuracy which all three algorithms have reached in 10 seeds. The results show FedLog needs less rounds to converge than LG-FedAvg and FedProto, thus saving more communication cost despite the cost per round is the same.
4.3 Celeba
We compare FedLog to LG-FedAvg and FedProto on large dataset Celeba [38] preprocessed by LEAF [39]. We sampled 70838 images from 2360 clients. Each client represents a celebrity, and the local data are images of the same person. Sample inputs are shown in the left half of Fig. 2. The task is to classify if the celebrity is smiling. MobileNetV2 [40] is implemented as the classifier model for all 3 algorithms. To test how these algorithms perform under the situation where only a few communication rounds are available, we run every algorithm to optimum locally before global aggregation. The results are shown in the right half of Fig. 2. FedLog performs statistically significantly better than FedProto (71.19$\pm$0.45 v.s. 69.98$\pm$0.16, $p=0.03$) and LG-FedAvg (66.84$\pm$0.19, $p=0.001$).
4.4 Flexible Architecture
We simulate the situation where clients have different computational resources on CIFAR100. We randomly select half of the clients and assign them a smaller convolutional neural network, where the second last fully connected layer is removed. Different clients start from different local body initialization, but the global head is unified. Other experiment details are similar to Sec. 4.2. We compare FedLog to LG-FedAvg 1 and FedProto. As shown in the left graph of Fig. 3, FedLog converges faster than LG-FedAvg and FedProto. The accuracy of FedLog is also statistically significantly higher than the accuracy of FedProto (55.31$\pm$0.15 v.s. 54.74$\pm$0.13, $p=0.002$) and LG-FedAvg (54.03$\pm$0.08, $p=0.001$).
Figure 2: Left: sample Celeba input. Right: model accuracy. Colored area shows mean $\pm$ standard deviation of the accuracy.
Figure 3: Left: flexible architecture results. Right: differential privacy trade-off. Colored area shows mean $\pm$ standard deviation of the accuracy.
4.5 Differential Privacy
We conduct experiments to show the trade-off between privacy budget $\epsilon$ and the accuracy of FedLog on CIFAR10. Experiment details are similar to Sec. 4.2, except we add an activation function to clip the extracted features to $b=2$, and $\delta=0.01$. As shown in the right graph of Fig. 3, the accuracy of differentially private FedLog quickly grows back to optimum when $\epsilon\geq 0.1$. This is a very strong privacy budget, and the results show FedLog performs well without sacrificing clients’ privacy.
5 Conclusion
We proposed FedLog that shares local data summaries instead of model parameters. FedLog assumes an exponential family model on local data, and learns a global linear separator with the summation of sufficient statistics. FedLog can learn universal local representations without sharing the bodies. Experiments show statistically significant improvements compared to prior arts, with the least communication cost. It is also effective with flexible architectures and formal DP guarantees. One limitation of FedLog is the assumption that the embedding of the data (just prior to the classification head) follows a mixture of Gaussians. Relaxing this assumption in future work could further improve the performance of the approach.
Acknowledgments and Disclosure of Funding
Resources used in preparing this research at the University of Waterloo were provided by Huawei Canada, the province of Ontario and the government of Canada through CIFAR and companies sponsoring the Vector Institute.
Appendix A Differential Privacy
It has been shown that FL algorithms that share large model parameters do not prevent privacy attacks through weight manipulation, GAN-based reconstruction, and large model memorization effects [41, 42].In our case, We argue that FedLog, which shares summations of sufficient statistics only, avoid these pitfalls closely related to model parameter sharing. Since “addition” is a non-invertible function, malicious attackers cannot recover features of individual data points. Since the local architecture and weights of the feature extractor is never shared in anyway, malicious attackers should not be able to reconstruct the original inputs, even if they are given the features of individual data points. We acknowledge that this argument is merely intuitive, and sharing data summaries could pose other risks of privacy leakage. To further guarantee users’ privacy formally, we now extend FedLog to be differentially private.
($\epsilon,\delta$)-DP protects clients’ privacy by adding noise to the shared information so that the adversaries cannot effectively tell if any record is included in the dataset (controlled by $\epsilon>0$) at most times (controlled by $10\leq\delta<$ 1) [43].
Definition A.1.
A mechanism $M_{DP}$ satisfies $(\epsilon$, $\delta)$-DP if for any two datasets $\mathcal{D},\mathcal{D}^{\prime}$ that differ by only one record (i.e. $|\mathcal{D}-\mathcal{D}^{\prime}\cup\mathcal{D}^{\prime}-\mathcal{D}|=1$), and for any possible output $O\in Range(M_{DP})$,
$\Pr_{O\sim M_{DP}(\mathcal{D})}\left[\log(\frac{\Pr[M_{DP}(\mathcal{D})=O]}{\Pr[M_{DP}(\mathcal{D}^{\prime})=O]})>\epsilon\right]<\delta$
Intuitively, $(\epsilon,\delta)$-DP guarantees that the inner log ratio, considered as the information loss leaked to the adversaries, is bounded by the privacy budget $\epsilon$ with probability $\delta$. Usually, $\epsilon\leq 1$ is viewed as a strong protection, while $\epsilon\geq 10$ does not protect much. The magnitude of noise needed is usually determined by $\epsilon,\delta$ and the sensitivity of the function $f$, of which the results ($f(\mathcal{D})$ the revealed information) need protection.
Definition A.2.
The $L_{p}$ sensitivity of any function $f:\mathcal{D}\xrightarrow{}\mathbb{R}^{n}$ is $L_{p}(f)=\max_{\mathcal{D},\mathcal{D}^{\prime}}||f(\mathcal{D})-f(\mathcal{D}^{\prime})||_{p}$. $\mathcal{D}$ and $\mathcal{D}^{\prime}$ differ by only one record.
A commonly used mechanism is to add Gaussian noise to $f(\mathcal{D})$:
Theorem A.3 (e.g. [43]).
A Gaussian Mechanism $\mathcal{GM}$ is $(\epsilon,\delta)$-DP with information revealing function $f$: $\mathcal{GM}(\mathcal{D})=f(\mathcal{D})+\mathcal{N}({\bf 0},\sigma^{2}{\bf I})$, where $\sigma={L_{2}(f)}\sqrt{2\ln(1.25/\delta)}/{\epsilon}$.
In FedLog, the only private information shared by clients is the summation of statistics ${\bf T}(\boldsymbol{\Phi}_{c},{\bf y}_{c})$, a vector of size $n_{class}*m$. Unfortunately, $L_{2}({\bf T})$ is unbounded for standard deep neural networks $\tilde{\boldsymbol{\theta}}_{c}$, since the output features are usually unbounded. We need to clip the absolute values of the features to $b$, by simply adding an activation function to the last layer of the feature extractor
$g(x):=\begin{cases}b,&\text{if }x>b\\ -b,&\text{if }x<-b\\ x,&\text{otherwise}\end{cases}$ (16)
Then we can calculate the $L_{2}$ sensitivity as follows
$\begin{split}L_{2}({\bf T})&=\max_{\mathcal{D},\mathcal{D}^{\prime}}||{\bf T}(\mathcal{D})-{\bf T}(\mathcal{D}^{\prime})||_{2}\\ &=\max_{\boldsymbol{\phi},y}||{\bf T}(\boldsymbol{\phi},y)||_{2}\\ &=\sqrt{1+(m-1)*b^{2}}\end{split}$ (17)
Appendix B Experiment Details
Communication cost, flexible architecture, and differential privacy experiments are run on 1 NVIDIA T4 GPU with 16GB RAM. Celeba experiments are run on 1 NVIDIA A40 GPU with 48GB RAM. Training data are normalized and randomly cropped and flipped. The architecture of CNNs used are listed in Table A.1. Some important hyperparameters are listed in Table A.2, A.3, and A.4. Most hyperparameters follow the experiment setting reported in LG-FedAvg. We make our code public in the supplementary materials, where further details can be found.
Appendix C Licences
Yann LeCun and Corinna Cortes hold the copyright of MNIST dataset, which is a derivative work from original NIST datasets. MNIST dataset is made available under the terms of the Creative Commons Attribution-Share Alike 3.0 license.
The CIFAR-10 and CIFAR-100 are labeled subsets of the 80 million tiny images dataset. They were collected by Alex Krizhevsky, Vinod Nair, and Geoffrey Hinton, made public at the CIFAR homepage.
The CelebA dataset is available for non-commercial research purposes only. See CelebA homepage for the full agreement.
MNIST CIFAR10 CIFAR100
nn.Conv2d(1, 10, kernel_size=5) nn.Conv2d(3, 6, kernel_size=5) nn.Conv2d(3, 6, kernel_size=5)
F.max_pool2d(kernel_size=2) nn.MaxPool2d(2,2) nn.MaxPool2d(2,2)
nn.Conv2d(10, 20, kernel_size=5) nn.Conv2d(6, 16, kernel_size=5) nn.Conv2d(6, 16, kernel_size=5)
F.max_pool2d(kernel_size=2) nn.MaxPool2d(2,2) nn.MaxPool2d(2,2)
nn.Linear(320, 50) nn.Linear(400, 120) nn.Linear(400, 120)
nn.Linear(50, 10) nn.Linear(120, 100) nn.Linear(120, 100)
nn.Linear(100, 10) nn.Linear(100, 100)
Table A.1: CNN architectures used in the communication cost experiment. Dropout layers and ReLu activation functions are omitted.
Table A.2: Hyperparameters used in communication cost experiments for MNIST.
Algorithm Hyperparameter Value
FedLog optimizer Adam
body learning rate 0.001
head learning rate 0.01
batch size 10
local epochs 5
FedAvg optimizer Adam
learning rate 0.001
batch size 10
local epochs 5
LG-FedAvg 1 # global layers 1
optimizer Adam
learning rate 0.001
batch size 10
local epochs 5
LG-FedAvg 2 # global layers 2
optimizer Adam
learning rate 0.001
batch size 10
local epochs 5
FedPer optimizer Adam
learning rate 0.001
batch size 10
local epochs 5
FedRep optimizer Adam
learning rate 0.001
batch size 10
body epochs 5
head epochs 10
CS-FL optimizer Adam
phase 1 learning rate 0.001
phase 2 learning rate 0.001
sparcity 0.005
dimension reduction 0.1
batch size 10
local epochs 5
Table A.3: Hyperparameters used in communication cost experiments for CIFAR10.
Algorithm Hyperparameter Value
FedLog optimizer Adam
body learning rate 0.0005
head learning rate 0.01
batch size 50
local epochs 1
FedAvg optimizer Adam
learning rate 0.0005
batch size 50
local epochs 1
LG-FedAvg 1 # global layers 1
optimizer Adam
learning rate 0.0005
batch size 50
local epochs 1
LG-FedAvg 2 # global layers 2
optimizer Adam
learning rate 0.0005
batch size 50
local epochs 1
FedPer optimizer Adam
learning rate 0.0005
batch size 50
local epochs 1
FedRep optimizer Adam
learning rate 0.0005
batch size 50
body epochs 1
head epochs 10
CS-FL optimizer Adam
phase 1 learning rate 0.001
phase 2 learning rate 0.01
sparcity 0.0005
dimension reduction 0.2
batch size 10
local epochs 1
Table A.4: Hyperparameters used in communication cost experiments for CIFAR100.
Algorithm Hyperparameter Value
FedLog optimizer Adam
body learning rate 0.0005
head learning rate 0.01
batch size 50
local epochs 3
FedAvg optimizer Adam
learning rate 0.0005
batch size 50
local epochs 3
LG-FedAvg 1 # global layers 1
optimizer Adam
learning rate 0.0005
batch size 50
local epochs 3
LG-FedAvg 2 # global layers 2
optimizer Adam
learning rate 0.0005
batch size 50
local epochs 3
FedPer optimizer Adam
learning rate 0.0005
batch size 50
local epochs 3
FedRep optimizer Adam
learning rate 0.0005
batch size 50
body epochs 3
head epochs 3
CS-FL optimizer Adam
phase 1 learning rate 0.001
phase 2 learning rate 0.01
sparcity 0.0005
dimension reduction 0.1
batch size 10
local epochs 1
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We achieved the smallest message size and fastest convergence without losing accuracy as shown in the experiment section. We illustrated the trade-off of the DP version.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See section 3.2 and the conclusion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Full proof of theorem 3.1 is in Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Experiment details are described along with the results and in Appendix B.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: See supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Full details are provided with the public code.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: One-tailed Wilcoxon signed-rank tests are done for all experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Appendix B.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification:
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: This paper presents work whose goal is to advance the field of Machine Learning. There could be potential societal consequences of our work, none which we feel must be specifically highlighted here.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Datasets used for experiments are properly cited. See Appendix C for the licenses.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 Abstract
-In federated learning (FL), the common paradigm that FedAvg proposes and most algorithms follow is that clients train local models with their private data, and the model parameters are shared for central aggregation, mostly averaging.
-In this paradigm, the communication cost is often a challenge, as modern massive neural networks can contain millions to billions parameters.
-We suggest that clients do not share model parameters but local data summaries, to decrease the cost of sharing.
-We develop a new algorithm FedLog with Bayesian inference, which shares only sufficient statistics of local data.
-FedLog transmits messages as small as the last layer of the original model.
-We conducted comprehensive experiments to show we outperform other FL algorithms that aim at decreasing the communication cost.
-To provide formal privacy guarantees, we further extend FedLog with differential privacy and show the trade-off between privacy budget and accuracy.
-1 Introduction
-This work focuses on the design of techniques that are communication efficient and flexible for personalized federated learning (FL).
-The main challenge in FL is to enable communication between different clients and a server to facilitate the estimation of (personalized) aggregated models while keeping private the data in each client [1].
-Hence, instead of sharing the data into a central location, clients share parameters of their local model which are then aggregated.
-However, this may induce large communication costs when the models are large (millions to billions of parameters) [1] and most aggregation schemes require homogeneous architectures across clients [2, 3, 4, 5, 6].
-While the data never leaves the clients, the shared parameters may be thought as implicit data summaries since they indirectly capture information about the data.
-Given this perspective, one may ask: can we reduce communication costs by sharing small data summaries while preserving privacy and can clients utilize different model architectures?
-This paper describes a new algorithm called FedLog that shares data summaries with much smaller dimensionality than model parameterization.
-Furthermore, clients may use different model architectures to compute those summaries.
-Since the parameters of those local models are never shared, they do not need to use the same architecture.
-Data summaries are aggregated in a central server that trains a global prediction head by Bayesian inference.
-More specifically, it treats the data summaries as sufficient statistics in an exponential family model and returns to the clients the canonical parameters of this exponential family model.
-While all clients use this same head for prediction, their locally trained models for data summarization enable them to personalize their predictions.
-Differential privacy (DP) is incorporated to mitigate privacy leakage via the data summaries.
-The paper makes the following contributions:
-  • FedLog: a new Federated Bayesian logistic regression algorithm, the model of which is carefully designed to provide a statistical interpretation for representation learning.
-  • Experiments demonstrating FedLog’s low communication cost (as small as 1.6% of the original model per round) and fast convergence under multiple situations.
+Federated representation learning (FRL) aims to learn personalized federated models with effective feature extraction from local data.
+FRL algorithms that share the majority of the model parameters face significant challenges with huge communication overhead.
+This overhead stems from the millions of neural network parameters and slow aggregation progress of the averaging heuristic.
+To reduce the overhead, we propose to share sufficient data summaries instead of raw model parameters.
+The data summaries encode minimal sufficient statistics of an exponential family, and Bayesian inference is utilized for global aggregation.
+It helps to reduce message sizes and communication frequency.
+To further ensure formal privacy guarantee, we extend it with differential privacy framework.
+Empirical results demonstrate high learning accuracy with low communication overhead of our method.
+Introduction
+Representation learning plays a crucial role in machine learning by effectively extracting features from raw data to facilitate downstream prediction in various fields (Liu, Shen, and Yang 2024).
+Combined with federated learning (FL), federated representation learning (FRL) aims to learn personalized federated models and diminish the impact of heterogeneous clients.
+FRL methods separate a whole neural network into two parts, body and head (Liang et al.
+2020; Collins et al.
+2021; Arivazhagan et al.
+2019).
+The body is a deep network that learns a compact feature representation from the raw data.
+The head is a shallow network with few layers that make predictions in the representation space.
+Personalization is achieved by localizing either the body or the head.
+Parameters of the rest of the model are shared with the server for global aggregation with the FedAvg heuristic (McMahan et al.
+2017).
+FRL utilizing body parameter sharing faces two significant challenges.
+1) Heavy communication overhead.
+Sharing body parameters induces huge communication cost per round due to millions of parameters of the deep learning model (Wen et al.
+2023) since the body contains the majority of the parameters.
+Furthermore, the averaging heuristic slows down the convergence with heterogeneous clients, resulting in more aggregation rounds.
+2) Rigid model architecture.
+Sharing body parameters requires the same body architecture among clients (McMahan et al.
+2017; Arivazhagan et al.
+2019; Collins et al.
+2021; Zhang et al.
+2024; Oh, Kim, and Yun 2021; Li, Li, and Varshney 2021).
+However, clients usually possess different amounts of computing resources.
+Thus, one body architecture may not be suitable for all devices.
+In this case, clients with limited resources cannot effectively participate.
+Even though some recent works consider sharing head parameters to reduce communication overhead, their communication efficiency is still limited due to slow averaging (Karimireddy et al.
+2020).
+To tackle the above challenges, our motivation is that a succinct data sharing protocol should minimize the bandwidth usage and does not depend on specific model architecture.
+Note that sharing model parameters in original FRL can be interpreted as sharing implicit client data summaries, since local model parameters capture the information from the input which can recover the raw data (Mothukuri et al.
+2021).
+However, there is no guarantee that the information reflected by these parameters is sufficient or necessary to infer the global model.
+Instead, we can consider sharing concise, but sufficient client data summaries.
+Sharing sufficient data summaries (also known as sufficient statistics (Jordan 2009a)) of clients offers two benefits.
+1) sufficient statistics maintain small data sizes which reduce the bandwidth usage.
+2) sufficient statistics are model independent, which allows heterogenous model deployment.
+Based on this idea, we propose Federated Bayesian Logistic Regression (FedLog), a new FRL strategy.
+We consider the case that the body of the model is localized and the head is updated by the server.
+FedLog acquires sufficient statistics from the client data encoded by the body with an exponential family distribution.
+The sufficient statistics from each client are then sent to the server to determine the optimal head parameters by maximizing the posterior with Bayesian inference.
+The theoretical property of sufficient statistics ensures that sufficient information is captured with fixed size to infer global model parameters.
+Also, due to the Bayesian inference step, the number of communication rounds is reduced thus improving communication efficiency.
+Note that since the model body is not shared and summation of sufficient statistics is non-invertible, FedLog also avoids potential privacy attacks through weight manipulation, GAN-based reconstruction, or large model memorization effects (Boenisch et al.
+2021; Mothukuri et al.
+2021).
+However, to ensure a formal privacy guarantee, we further incorporate differential privacy framework to mitigate privacy leakage.
+In summary, the paper makes the following contributions:
+  • FedLog: a new FRL algorithm, the model of which is carefully designed to provide a statistical interpretation for representation learning.
+  • Experiments demonstrating FedLog’s low communication cost (as small as $0.09\%$ of FedAvg) and fast convergence under multiple scenarios.
   • Incorporation of DP and demonstration of favorable trade-off between privacy and utility.
-2 Background
-2.1 Federated Learning and Personalized FL
-In Federated Learning (FL), there is always a collection of clients $c\in S$ who wish to collaborate, and each client holds their own data $\mathcal{D}_{c}=({\bf X}_{c},{\bf y}_{c})$ locally.
-We seek to train some machine learning model $\mathcal{M}_{\theta}$ with parameters $\boldsymbol{\theta}$ on these client data.
-Conventionally, we would centralize all the data $\mathcal{D}=\bigcup_{c\in S}\mathcal{D}_{c}$, and choose some algorithm $\mathcal{A}$ and parameter initialization $\boldsymbol{\theta}_{0}$ to learn the model:
-$\mathcal{M}_{\boldsymbol{\theta}}=\mathcal{A}(\mathcal{D},\boldsymbol{\theta}_{0})$
-This simple approach becomes infeasible if the clients cannot directly share their data due to privacy concerns.
+Related Works
+Federated Representation Learning
+In FL, there is a collection of clients $c\in S$ wishing to collaborate.
+Each client holds their own data $\mathcal{D}_{c}=({\bf X}_{c},{\bf y}_{c})$ locally.
+We seek to train some ML model with parameters $\boldsymbol{\theta}$ on these client data.
+Conventionally, we would centralize all the data $\mathcal{D}=\bigcup_{c\in S}\mathcal{D}_{c}$ and learn the model with $\mathcal{D}$.
+This approach becomes infeasible if the clients cannot share their data due to privacy concerns.
 FL intends to tackle this problem.
-A trusted central server is allowed to receive and send perturbed matrices that only contain limited information about raw client data, such as model parameters.
-Most FL algorithms follow the following paradigm.
-Such algorithms iterate between local optimization and global aggregation.
-During local optimization at $t^{th}$ iteration, all the clients (or a proportion of the clients in some cases) start with the global consensus parameters $\theta^{t}$, and run the algorithm $\mathcal{A}$ to update the local models: $\mathcal{M}_{\boldsymbol{\theta}^{t+1}_{c}}=\mathcal{A}(\mathcal{D}_{c},\boldsymbol{\theta}^{t})$.
-Note the local models could have other local parameters $\tilde{\boldsymbol{\theta}}$ than $\boldsymbol{\theta}$, but is omitted in the notation for simplicity.
-Then, clients potentially compress their updated model parameters, and send them instead of data to the central server so that it can be aggregated into the global consensus parameters for next iteration $\boldsymbol{\theta}^{t+1}$.
-In this paradigm, one challenge is how to aggregate model parameters so that the resulting global consensus $\boldsymbol{\theta}^{t+1}$ is a better approximation to the centralization version.
-This is especially challenging when the number of local training epochs $>1$ [7], due to non-linear loss functions and predictors, thus many FL algorithms simply resort to the averaging heuristic [2, 3, 4, 5, 6].
-For example, the very first FL algorithm, FedAvg [2], let the clients send all model parameters to the server without compressing, and their weighted average becomes the new global consensus.
-Another challenge is that clients often have different data distribution $\Pr({\bf X}),\Pr({\bf y})$ or even $\Pr({\bf y}|{\bf X})$.
-This is referred to as non-i.i.d.
-clients, and it could harm the global aggregation in terms of convergence rate and model utility [8], especially when the goal is to learn a global model.
-Thus, personalized FL is introduced, where a global model is not mandatory, but each client could have their own model that best fit their data distribution.
-See [9] for more motivation and algorithms.
-2.2 Exponential Family
+A trusted central server is allowed to receive and send perturbed matrices that only contain limited information about raw data, such as model parameters.
+One challenge of FL is how to aggregate model parameters learnt locally so that the resulting global consensus $\boldsymbol{\theta}^{t+1}$ is a better approximation to the centralization version than the last round $\boldsymbol{\theta}^{t}$ This is especially challenging when the number of local training epochs $>1$ (Karimireddy et al.
+2020), due to non-linear loss functions and predictors.
+Thus, many FL algorithms simply resort to the averaging heuristic (McMahan et al.
+2017; Liang et al.
+2020; Arivazhagan et al.
+2019; Collins et al.
+2021; Achituve et al.
+2021).
+In FL, clients often have different data distribution $\Pr_{c}({\bf X}),\Pr_{c}({\bf y})$ or even $\Pr_{c}({\bf y}|{\bf X})$.
+This is referred to as heterogeneous or non-i.i.d.
+clients.
+The averaging heuristic can drastically harm the global aggregation in terms of convergence rate and model utility with non-i.i.d.
+clients (Li et al.
+2019).
+Thus, personalized FL (PFL) is introduced, where a global model is not mandatory, but each client could have their own model that best fits their data distribution (Tan et al.
+2022a).
+Nevertheless, most algorithms still utilize the averaging heuristic for aggregation.
+One line of FRL works approach PFL by localizing either the body (LG-FedAvg (Liang et al.
+2020)) or the head (FedPer (Arivazhagan et al.
+2019), FedRep (Collins et al.
+2021)) of the client models, and shares the rest with the server for averaging.
+FedProto (Tan et al.
+2022b) localizes the whole model, but averages feature representations by class and forces local models to learn similar representations.
+FedLog can also be interpreted as a representation learning algorithm, where we learn local representations by all the layers except the last one.
+However, unlike previous works that share model parameters and heuristically average them, we share sufficient statistics and update the global head with Bayesian inference.
+Another line of works (CCVR (Luo et al.
+2021), FedPFT (Beitollahi et al.
+2024)) fit Gaussian Mixture Models (GMM) to local features extracted by a globally uniformed body, and shares the natural parameters.
+The server then draws virtual features from the GMM and trains a global model.
+It is notable that these algorithms rely on pretrained FedAvg or foundation models to unify the body.
+Although FedLog also fits an exponential family distribution to local features, we do not need such pretrained models since our bodies are trained locally with any architecture.
+Also, we work with the canonical parameters and do not need to draw virtual samples from the learnt distribution.
+Other related works include communication efficient FL and Bayesian FL. See Appendix A for more details.
+Method
+Exponential Family
+We start by introducing the definition of exponential family.
+Definition 1.
 Exponential family refers to a set of probability distributions of the following canonical form.
 $\Pr({\bf x}|\boldsymbol{\eta})=h({\bf x})\exp(\boldsymbol{\eta}^{\top}{\bf T}({\bf x})-A(\boldsymbol{\eta}))$ (1)
+where $h({\bf x}):\mathbb{R}^{p}\rightarrow\mathbb{R}_{\geq 0}$, ${\bf T}({\bf x}):\mathbb{R}^{p}\rightarrow\mathbb{R}^{d}$, $A(\boldsymbol{\eta}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ are known functions.
 […]
-2.3 Conjugate Priors
-Suppose we wish to model the distribution of some data ${\bf x}$ by learning some parameters ${\boldsymbol{\eta}}$, and denote the likelihood induced by the model $\Pr({\bf x}|\boldsymbol{\eta})$.
-With Bayesian inference, $\boldsymbol{\eta}$ is often treated as random variables, and a prior distribution $\Pr(\boldsymbol{\eta})$ can be specified to incorporate prior knowledge.
-Then, by Bayes’ Theorem the posterior distribution is:
-[…]
-The integral in Eq.
-2 and thus the posterior $\Pr(\boldsymbol{\eta}|{\bf x})$ may not be tractable for arbitrary priors $\Pr(\boldsymbol{\eta})$.
-A convenient choice that guarantees analytical solutions is the conjugate prior.
-Given a likelihood $\Pr({\bf x}|\boldsymbol{\eta})$, a prior is called its conjugate prior if $\Pr(\boldsymbol{\eta})$ and $\Pr({\bf x}|\boldsymbol{\eta})$ follow the same distribution family.
-Specifically, if the likelihood is an exponential family, it has known conjugate priors [11].
-Likelihood: $\Pr({\bf x}|\boldsymbol{\eta})=h({\bf x})\exp(\boldsymbol{\eta}^{\top}{\bf T}({\bf x})-A(\boldsymbol{\eta}))$ (3)
-Prior: $\Pr(\boldsymbol{\eta};\boldsymbol{\chi},\nu)=f(\boldsymbol{\chi},\nu)\exp(\boldsymbol{\eta}^{\top}\boldsymbol{\chi}-{\nu}A(\boldsymbol{\eta}))$ (4)
-Posterior: $\Pr(\boldsymbol{\eta}|{\bf x})=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+{\bf T}({\bf x}),\nu+1)$ (5)
-[…]
-2.4 Related Work
-Representation learning.
-The most related to our works are federated representation learning methods.
-These approaches separate a whole neural network into two parts, body and head.
-They interpret the body as a deep neural network that learns a compact feature representation from raw data, and the head as a few shallow layers that map the features to the prediction.
-LG-FedAvg [3] chooses to localize the body and learn local feature representations, but averages the head for global information sharing.
-On the contrary, FedPer [4] localizes the head but averages the body to learn a universal representation.
-Later, FedRep [5] also localizes the head, but instead of jointly updating the body and the head locally, they first update the local head to convergence, and then fix the head and update the body.
-FedProto [12] localizes the whole model, but averages feature representations by class and forces local models to learn similar representations.
-FedLog can also be interpreted as a representation learning algorithm, where we learn local representations by all the layers except the last one.
-However, unlike previous works that share model parameters and heuristically average them, we share data summaries and update the global head with Bayesian inference.
-Another line of works (CCVR [13], FedPFT [14]) fit Gaussian Mixture Models (GMM) to local features extracted by a globally uniformed body, and shares the natural parameters.
-The server then draws virtual features from the GMM and train a global model.
-It is notable that these algorithms rely on pretrained FedAvg or foundation models to unify the body.
-Although FedLog also fits a exponential family distribution to local features, we do not need such pretrained models since our bodies are trained locally with any architecture.
-Also, we work with the canonical parameters (which result in smaller messages) and do not need to draw virtual samples from the learnt distribution.
-Communication efficient FL.
-In addition to localizing part of the original model, which decreases the message size straightforwardly, other lines of work that aim at communication efficient FL include optimization algorithms, client selection, and model compression [1].
-First, it has been shown that local training epochs could impact the rounds of global communication needed [2].
-Based on this observation, researchers proposed different local optimization methods to reduce the communication rounds [15, 16, 17].
-On the other hand, instead of updating all clients or randomly selecting a subset of clients, researchers proposed various algorithms that accelerate the global learning process by carefully selecting clients that contribute more to the global model or are more efficient in terms of uploading time [18, 19, 20, 21].
-Additionally, researchers proposed algorithms that transmit smaller messages than the whole model by reducing and compressing the parameters [22, 23, 24].
-However, most such algorithms are efficient at the cost of model accuracy [25].
-Bayesian FL.
-Other Bayesian models have been explored to represent distributions over models and predictions in FL.
-The challenge is in the aggregation of the local posteriors.
-Various techniques have been proposed including personalized GPs [6], posterior averaging [26], online Laplace approximation [27], Thompson sampling [28], MCMC sampling [29].
-There is not much in common between these techniques with our approach despite the use of Bayes theorem since prior Bayesian techniques tend to emphasize on calibration, approximating the posterior, or even different tasks.
-3 Federated Bayesian Logistic Regression
-In this section, we first detail our model construction and algorithm. Then, we show how to incorporate differential privacy to prevent potential privacy leakage.
-3.1 FedLog
-Notations.
+Many well-known distributions are included in the exponential family, such as Gaussian, binomial, Poisson, and Bernoulli distribution.
+We can always transform a distribution represented by natural parameters into its canonical form defined above.
+For example, a binomial distribution has the following p.d.f.:
+$\Pr(x|p)={n\choose x}p^{x}(1-p)^{(n-x)},x\in\{0,1,\cdots,n\}$
+It can be rewritten as:
+$\Pr(x|\eta)={n\choose x}\exp({\eta}x-n\ln({1+e^{\eta}})),\eta=\ln\frac{p}{1-p}$
+Exponential family has a few desirable properties, which makes it a perfect candidate for information sharing in PFL.
+1) The sufficient statistic ${\bf T}({\bf x})$ captures all the information of ${\bf x}$ that can be used to infer $\boldsymbol{\eta}$ (i.e., ${\bf x}{\perp\!\!\!\perp}\boldsymbol{\eta}|{\bf T(x)}$, conditional independence) (Jordan 2009a).
+2) If ${\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{n}$ are i.i.d.
+samples from $\Pr({\bf x}|{\boldsymbol{\eta}})$, the summation of sufficient statistics $\sum_{i=1}^{n}{\bf T}({\bf x}_{i})$ is a complete statistic for $\boldsymbol{\eta}$.
+It contains only information about $\boldsymbol{\eta}$, without ancillary information (Casella and Berger 2015).
+3) It is the only parametric distribution family with sufficient statistics of fixed size that does not grow with the sample size (Koopman 1936).
+4) It has known conjugate priors (Jordan 2009b), a crucial property for Bayesian inference as we will elaborate later.
+Formal definitions of these properties are in Appendix B.
+Model Construction
+We detail our model construction and algorithm with the following notation.
+Notation.
 Let $\tilde{\boldsymbol{\theta}}_{c}$ denote the parameters of a local neural network of client $c$.
+This network serves as a local body.
 Let $f_{\tilde{\boldsymbol{\theta}}_{c}}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{m}$ be the function of the local neural network and $\boldsymbol{\Phi}_{c}=f_{\tilde{\boldsymbol{\theta}}_{c}}({\bf X}_{c})$ be the $m$ local features extracted from the local input.
-For convenience, we designate one of the features to be always $1$.
+For convenience, we designate the first feature to be always $1$.
 Let $n_{class}$ denote the total number of classes in a classification task, and $n_{c}$ denote the size of the local dataset.
@@
 one-hot vector form) of label $i\in\{1,2,\cdots,n_{class}\}$, and $\otimes:\mathbb{R}^{m}\times\mathbb{R}^{n_{class}}\rightarrow\mathbb{R}^{m*n_{class}}$ be the Kronecker product.
-We assume the joint probability of each data point in $(\boldsymbol{\Phi}_{c},{\bf y}_{c})$, denoted as $\Pr(\boldsymbol{\phi},y)$, is an exponential family with canonical parameters $\boldsymbol{\eta}\in\mathbb{R}^{m*n_{class}}$ as follows.
-sufficient statistics: ${\bf T}(\boldsymbol{\phi},y):=\boldsymbol{\phi}\otimes{\bf e}_{y}$ (6)
-base measure: $h(\boldsymbol{\phi},y):=\frac{\exp\left(-\sum_{i=1}^{m}\boldsymbol{\phi}_{i}^{2}\right)}{\sqrt{\pi^{m}}}$ (7)
+We assume the joint probability of each data point in $(\boldsymbol{\Phi}_{c},{\bf y}_{c})$, denoted as $\Pr(\boldsymbol{\phi},y)$, is an exponential family with canonical parameters $\boldsymbol{\eta}\in\mathbb{R}^{m*n_{class}}$.
+We design the sufficient statistics ${\bf T}(\boldsymbol{\phi},y)$ and the base measure $h(\boldsymbol{\phi},y)$ to have the following form.
+${\bf T}(\boldsymbol{\phi},y)$ $:=\boldsymbol{\phi}\otimes{\bf e}_{y}$ (2)
+$h(\boldsymbol{\phi},y)$ $:=\frac{\exp\left(-\sum_{i=1}^{m}\boldsymbol{\phi}_{i}^{2}\right)}{\sqrt{\pi^{m}}}$ (3)
 where $\boldsymbol{\phi}_{i}\in\mathbb{R}$ denotes the $i^{th}$ entry of $\boldsymbol{\phi}$.
-One advantage with this specific choice of $\bf T$ and $h$ is that $A$ has an explicit expression.
+One advantage with this specific choice of $\bf T$ and $h$ lies in the resulting conditional likelihood $\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$.
 Let $\boldsymbol{\eta}_{y}\in\mathbb{R}^{m}$ denote the $((y-1)*m)^{th}$ to $(y*m)^{th}$ entries of $\boldsymbol{\eta}$, $y\in\{1,2,\cdots,n_{class}\}$.
-Let $\boldsymbol{\eta}_{y,i}\in\mathbb{R}$ denote the $i^{th}$ entry of $\boldsymbol{\eta}_{y}$.
 Then,
-[…]
-We can write out the likelihood, conditional likelihood, prior, and posterior as follows.
-$\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})$ $=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}-\boldsymbol{\phi}^{\top}\boldsymbol{\phi})}{\sqrt{\pi^{m}}\sum_{y^{\prime}=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\eta}_{y^{\prime}}/4)}$ (9)
-$\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$ $=\frac{\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})}{\Pr(\boldsymbol{\phi}|\boldsymbol{\eta})}=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi})}{\sum_{y^{\prime}=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\phi})}$ (10)
-$\Pr(\boldsymbol{\eta};\boldsymbol{\chi},\nu)$ $=f(\boldsymbol{\chi},\nu)\frac{\exp(\boldsymbol{\eta}^{\top}\boldsymbol{\chi})}{(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\eta}_{y^{\prime}}/4))^{\nu}}$ (11)
-$\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},y)$ $=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\phi}\otimes{\bf e}_{y},\nu+1)$ (12)
-Unfortunately, the normalizing constant function $f(\boldsymbol{\chi},\nu)$, the marginal likelihood $\Pr(y|{\boldsymbol{\phi}})$, and the predictive posterior $\Pr(y_{*}|\boldsymbol{\phi}_{*},\boldsymbol{\phi},y)$ are still intractable.
-However, another advantage of our choice of ${\bf T}$ and $h$ is shown by Eq.
-10: the conditional likelihood is exactly the softmax function over $\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}$.
+$\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})$ $=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}-\boldsymbol{\phi}^{\top}\boldsymbol{\phi}-A({\boldsymbol{\eta}}))}{\sqrt{\pi^{m}}}$ (4)
+$\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$ $=\frac{\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})}{\Pr(\boldsymbol{\phi}|\boldsymbol{\eta})}=\frac{\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})}{\sum_{y^{\prime}=1}^{n_{class}}\Pr(\boldsymbol{\phi},y^{\prime}|\boldsymbol{\eta})}$ (5)
+$=\frac{\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi})}{\sum_{y^{\prime}=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y^{\prime}}^{\top}\boldsymbol{\phi})}$ (6)
+Eq.
+6 is exactly the softmax function over $\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\phi}$.
 This means we can take any deep neural network that extracts $m$ features, and append $\boldsymbol{\eta}$ as the last linear layer that maps the features to $n_{class}$ logits.
 Then, this composed neural network serves as a stand-alone classifier that computes the conditional likelihood $\Pr(y|{\boldsymbol{\phi}},\boldsymbol{\eta})$.
+However, Eq.
+6 cannot be directly maximized at the server, since it requires knowledge of uncompressed representation-label pairs.
+Instead, we utilize Bayesian inference to optimize $\boldsymbol{\eta}$.
+With Bayesian inference, $\boldsymbol{\eta}$ is treated as a random variable.
+A prior distribution $\Pr(\boldsymbol{\eta})$ can be specified to incorporate prior knowledge.
+Then, by Bayes’ Theorem, the posterior distribution is:
+$\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},y)$ $=\frac{\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})\Pr(\boldsymbol{\eta})}{\Pr(\boldsymbol{\phi},y)}$ (7)
+[…]
+The integral in Eq.
+8 and thus the posterior $\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},y)$ may not be tractable for arbitrary priors $\Pr(\boldsymbol{\eta})$.
+A convenient choice that guarantees analytical solutions is the conjugate prior.
+Given a likelihood $\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})$, a prior is called its conjugate prior if $\Pr(\boldsymbol{\eta})$ and $\Pr(\boldsymbol{\phi},y|\boldsymbol{\eta})$ follow the same distribution family.
+Specifically, if the likelihood is an exponential family, it has known conjugate priors (Jordan 2009b).
+Prior: $\Pr(\boldsymbol{\eta};\boldsymbol{\chi},\nu)=f(\boldsymbol{\chi},\nu)\exp(\boldsymbol{\eta}^{\top}\boldsymbol{\chi}-{\nu}A(\boldsymbol{\eta}))$ (9)
+Posterior: $\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},y)=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+{\bf T}(\boldsymbol{\phi},y),\nu+1)$ (10)
+[…]
+Another advantage of our model design is that $A$ has an explicit expression. Let $\boldsymbol{\eta}_{y,i}\in\mathbb{R}$ denote the $i^{th}$ entry of $\boldsymbol{\eta}_{y}$.
+[…]
+Due to this analytical solution, we can directly optimize Eq. 10 without further approximations.
+FedLog
+Based on the above model, we propose our new algorithm FedLog, summarized in Algo.
+1.
+At the beginning, the server initializes $\boldsymbol{\eta}$ (the global head) randomly.
+The clients initialize $\tilde{\boldsymbol{\theta}}_{c}$ (the local bodies) either completely randomly, or with the same random seed sent by the server to unify the initialization.
+Note $\tilde{\boldsymbol{\theta}}_{c}$ is not part of our exponential family assumption, thus we do not require them to have the same shape or architecture amongst different clients.
+Parameters that we need to optimize are essentially $\tilde{\boldsymbol{\theta}}_{c}$ and $\boldsymbol{\eta}$, which can be done by maximizing $\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$ and $\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},{y})$ in turns iteratively, similarly to the expectation-maximization algorithm.
+Concretely, all the clients $c\in S$ first fix the global head $\boldsymbol{\eta}$, and update their local bodies $\tilde{\boldsymbol{\theta}}_{c}$ with gradient descent.
+We derive the loss function as:
+$\begin{split}\mathcal{L}_{c}&=-\sum_{i=1}^{n_{c}}\ln\Pr({\bf y}_{c,i}|\boldsymbol{\Phi}_{c,i},\boldsymbol{\eta})\\ &=-\sum_{i=1}^{n_{c}}\ln\frac{\exp(\boldsymbol{\eta}_{y_{c,i}}^{\top}\boldsymbol{\Phi}_{c,i})}{\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\Phi}_{c,i})}\end{split}$ (12)
+where $(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})\in\mathbb{R}^{m}\times\mathbb{R}$ is the $i^{th}$ data point of $\boldsymbol{\Phi}_{c},{\bf y}_{c}$.
+This is exactly the cross entropy loss widely used in deep learning for classification tasks.
+Then, clients compute sufficient statistics of their local data $\sum_{i=1}^{n_{c}}{\bf T}(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})=\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}}$.
+They then send the summations and $n_{c}$ to the server for global head learning.
+As we have discussed with the introduction of exponential family, the sufficient statistics contain all information in the representations that could be used to infer $\boldsymbol{\eta}$ in our model.
+The server only needs to know the summation of all the sufficient statistics, since
+$\begin{split}&\Pr(\boldsymbol{\eta}|\boldsymbol{\Phi}_{c_{1}},{\bf y}_{c_{1}},\cdots,\boldsymbol{\Phi}_{c_{k}},{\bf y}_{c_{k}})\\ &=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},\nu+\sum_{c\in S}n_{c})\end{split}$ (13)
+can be trivially inferred from Eq.
+10.
+Note the size of the message sent by each client equals the size of the last linear layer.
+After receiving the sufficient statistics, the server computes $\boldsymbol{\Phi}=\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},n=\sum_{c\in S}n_{c}$ and updates the global head $\boldsymbol{\eta}$ by maximum a posteriori (MAP):
+$\begin{split}\boldsymbol{\eta}&=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)\\ &=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\frac{\exp\left(\boldsymbol{\eta}^{\top}(\boldsymbol{\chi}+\boldsymbol{\Phi})\right)}{(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\eta}_{y}/4))^{(\nu+n)}}\end{split}$ (14)
 Algorithm 1 FedLog (${\bf X}_{c},{\bf y}_{c}$: local data, $\tilde{\boldsymbol{\theta}}_{c}$ local body parameters, $\boldsymbol{\eta}$: global head parameters, $\boldsymbol{\chi}$: prior parameter, $\nu$: prior parameter, $\zeta$: local learning rate)
@@
      $\boldsymbol{\Phi}\leftarrow\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},n\leftarrow\sum_{c\in S}n_{c}$
-     $\boldsymbol{\eta}\leftarrow\operatorname*{arg\,max}_{\boldsymbol{\eta}}\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)$ (Eq.  15)
-Based on the current model, we propose a new algorithm named FedLog, as shown in Alg.
-1.
-At the beginning, the server initializes $\boldsymbol{\eta}$ (the global head) randomly, and the clients initializes $\tilde{\boldsymbol{\theta}}_{c}$ (the local bodies) either completely randomly, or with the same random seed sent by the server to unify the initialization.
-Note $\tilde{\boldsymbol{\theta}}_{c}$ is not part of our exponential family assumption, thus we do not require them to have the same shape or architecture amongst different clients.
-Parameters that we need to optimize are essentially $\tilde{\boldsymbol{\theta}}_{c}$ and $\boldsymbol{\eta}$, which can be done by maximizing $\Pr(y|\boldsymbol{\phi},\boldsymbol{\eta})$ and $\Pr(\boldsymbol{\eta}|\boldsymbol{\phi},{y})$ in turns iteratively, similarly to the expectation-maximization algorithm.
-Concretely, all the clients $c\in S$ first fix the global head $\boldsymbol{\eta}$, and update their local bodies $\tilde{\boldsymbol{\theta}}_{c}$ with gradient descent.
-We derive the loss function as:
-$\begin{split}\mathcal{L}_{c}&=-\sum_{i=1}^{n_{c}}\ln\Pr({\bf y}_{c,i}|\boldsymbol{\Phi}_{c,i},\boldsymbol{\eta})=-\sum_{i=1}^{n_{c}}\ln\frac{\exp(\boldsymbol{\eta}_{y_{c,i}}^{\top}\boldsymbol{\Phi}_{c,i})}{\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\Phi}_{c,i})}\end{split}$ (13)
-where $(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})\in\mathbb{R}^{m}\times\mathbb{R}$ is the $i^{th}$ data point of $\boldsymbol{\Phi}_{c},{\bf y}_{c}$.
-This is exactly the cross entropy loss widely used in deep learning for classification tasks.
-Then, clients compute sufficient statistics of their local data $\sum_{i=1}^{n_{c}}{\bf T}(\boldsymbol{\Phi}_{c,i},{\bf y}_{c,i})=\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}}$.
-They then send the summations and $n_{c}$ to the server for global head learning.
-As we have discussed in Sec.
-2.2, the sufficient statistics contain all the information in the raw dataset that could be used to infer $\boldsymbol{\eta}$ in our model.
-In this case, the server only needs to know the summation of all the sufficient statistics, since
-$\begin{split}&\Pr(\boldsymbol{\eta}|\boldsymbol{\Phi}_{c_{1}},{\bf y}_{c_{1}},\cdots,\boldsymbol{\Phi}_{c_{k}},{\bf y}_{c_{k}})=\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},\nu+\sum_{c\in S}n_{c})\end{split}$ (14)
-can be trivially inferred from Eq.
-12.
-Note the size of the message sent by each client equals the size of the last linear layer.
-After receiving the sufficient statistics, the server computes $\boldsymbol{\Phi}=\sum_{c\in S}\sum_{i=1}^{n_{c}}\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}},n=\sum_{c\in S}n_{c}$ and updates the global head $\boldsymbol{\eta}$ by maximum a posteriori (MAP):
-$\begin{split}\boldsymbol{\eta}&=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)=\operatorname*{arg\,max}_{\boldsymbol{\eta}}\ln\frac{\exp\left(\boldsymbol{\eta}^{\top}(\boldsymbol{\chi}+\boldsymbol{\Phi})\right)}{(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\eta}_{y}/4))^{(\nu+n)}}\end{split}$ (15)
+     $\boldsymbol{\eta}\leftarrow\operatorname*{arg\,max}_{\boldsymbol{\eta}}\Pr(\boldsymbol{\eta};\boldsymbol{\chi}+\boldsymbol{\Phi},\nu+n)$ (Eq.  14)
 Note Eq.
-15 is a convex optimization task, and we can easily compute its sole maximum by gradient descent.
+14 is a convex optimization task.
+We can easily compute its sole maximum by gradient descent with complexity $O(m*n_{class})$.
 Then, the server sends $\boldsymbol{\eta}$ back to all the clients and starts the next round of updates.
 The process is repeated until convergence.
-3.2 Interpret FedLog
+Figure 1: Synthetic experiments.
+Dots are data points or local representations: dark green: client 0 class 0; light green: client 1 class 0; Dark blue: client 0 class 1; light blue: client 1 class 1.
+Dashed lines are linear separators.
+Accuracy results are averaged over 6 seeds.
+Interpretation and FedLog-C
 In this section, we analyze our assumptions in more details and give some insights about how FedLog works.
 The first assumption we made is that $\forall c\in S$, the local data points $\boldsymbol{\Phi}_{c},{\bf y}_{c}$ are from the same exponential family distribution whose pdf is given by Eq.
-9.
+4.
 In other words, we assume that the local bodies $\tilde{\boldsymbol{\theta}}_{c}$ transform their input, of any form, to the same representation space.
-This may seem infeasible at the first glance, since we do not directly aggregate the local body parameters, and they may take any architecture.
+This may seem infeasible at first glance since we do not directly aggregate the local body parameters, and they may follow any architecture.
 However, note we fix the global head $\boldsymbol{\eta}$ during local updates, by which the local bodies are forced to learn a universal representation space.
-See Sec.
-4.1 for more details.
+See the synthetic experiment below for more details.
 This assumption allows principled discriminative training for the local bodies with cross entropy loss, but unavoidably leaves a generative model for learning the global head.
@@
 This means we essentially assumed a mixture of Gaussians for the local features $\boldsymbol{\phi}$.
-This is a mild assumption, since sufficiently expressive DNN encoders can map data to any distribution any closely.
+To mitigate the gap between the assumption and the actual feature distribution, we propose a variation FedLog-C.
+An auxiliary loss is added during local training to force the local bodies to learn Gaussian-like clusters.
+Let $\boldsymbol{\Phi}_{y}$ denote the $((y-1)*m)^{th}$ to $(y*m)^{th}$ entries of $\boldsymbol{\Phi}$.
+Let $\boldsymbol{\Phi}_{y,0}$ denote the first entry of $\boldsymbol{\Phi}_{y}$.
+Then $\overline{\boldsymbol{\Phi}_{y}}=\boldsymbol{\Phi}_{y}/\boldsymbol{\Phi}_{y,0}$ is the global mean representation of class $y$.
+Inspired by contrastive learning (Schroff, Kalenichenko, and Philbin 2015), we derive the new local loss as:
+$\begin{split}\mathcal{L}^{\prime}_{c}=\mathcal{L}_{c}&+\alpha\sum_{i=1}^{n_{c}}(\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}}-\overline{\boldsymbol{\Phi}_{{\bf y}_{c,i}}})^{2}/n_{c}\\ &-\beta\sum_{y^{\prime}\neq{\bf y}_{c,i}}\sum_{i=1}^{n_{c}}(\boldsymbol{\Phi}_{c,i}\otimes{\bf e}_{{\bf y}_{c,i}}-\overline{\boldsymbol{\Phi}_{y^{\prime}}})^{2}/n_{c}\end{split}$ (15)
+$\alpha\in\mathbb{R}_{\geq 0}$ controls how compact the clusters should be.
+$\beta\in\mathbb{R}_{\geq 0}$ controls the distance between different clusters.
+Since clients need to know $\boldsymbol{\Phi}$, the server simply broadcasts the aggregated statistic to all the clients.
+Clients can optimize the same $\boldsymbol{\eta}$ locally, preserving the same communication cost.
 The second assumption we made is the prior of the global head.
@@
 The prior then becomes $\Pr(\boldsymbol{\eta})\propto\exp(-A(\boldsymbol{\eta}))=(\sum_{y=1}^{n_{class}}\exp(\boldsymbol{\eta}_{y}^{\top}\boldsymbol{\eta}_{y}/4))^{-1}$.
-The pdf takes its maximum at $\boldsymbol{\eta}={\bf 0}$ and decreases quickly as the absolute value of entries of $\boldsymbol{\eta}_{y}$ grows larger.
-This is in analogy to the Lasso regularizer in the regression case, which prevents the model from learning coefficients with large absolute values that are more likely due to noise or over-fitting.
+The p.d.f.
+takes its maximum at $\boldsymbol{\eta}={\bf 0}$ and decreases quickly as the absolute value of entries of $\boldsymbol{\eta}_{y}$ grows larger.
+This is in analogy to the Lasso regularizer in the regression case, which prevents the model from learning coefficients with large absolute values due to noise or over-fitting.
 From the Bayesian view, FedLog can take any deep classifier, and make the last linear layer Bayesian.
-It essentially operates a Bayesian logistic regression model on the local representations, thus the name FedLog.
-We start from a generative assumption and achieve the cross entropy loss conditional likelihood usually assumed directly in Bayesian logistic regression.
-We obtained an analytical solution for the kernel of the posterior, which can be calculated easily by the summation of sufficient statistics, whose sizes are as small as the size of the last linear layer, and avoid the need to share separated data points.
+It essentially operates a Bayesian logistic regression model on the local representations.
+We start from a generative assumption and achieve the cross-entropy loss conditional likelihood usually assumed directly in Bayesian logistic regression.
+We obtained an analytical solution for the kernel of the posterior, which can be calculated easily by the summation of sufficient statistics.
+The shared statistic cannot be further compressed without losing information from the representations.
+We formalize this statement with the following theorem.
+Theorem 2.
+If $({\boldsymbol{\phi}_{1}},y_{1}),({\boldsymbol{\phi}_{2}},y_{2}),\cdots,({\boldsymbol{\phi}_{n}},y_{n})$ are i.i.d.
+samples from the exponential family defined with Eq.
+4, then ${\bf T}(({\boldsymbol{\phi}_{1}},y_{1}),\cdots,({\boldsymbol{\phi}_{n}},y_{n}))=\sum_{i=1}^{n}{\boldsymbol{\phi}_{i}}\otimes{\bf e}_{y_{i}}$ is a minimal sufficient statistic independent of every ancillary statistic.
+Proof.
+See Appendix B for the proof. ∎
 From the federated representation learning view, FedLog has a one-layer global head and a deep local body.
@@
 The server is only responsible for finding the best linear separator given the local representations.
-As we will show with the experiments, we can automatically learn universal local representations and converge faster than using the averaging heuristic.
-3.3 Differential Privacy
-FL can be combined with formal mechanisms such as differential privacy (DP) [30, 31] or secure multi-party computation (MPC) [32, 33, 34], to provide formal privacy guarantees.
+As we will show with the experiments, we can converge faster than using the averaging heuristic.
+Differential Privacy
+FL can be combined with formal mechanisms such as differential privacy (DP) (Wei et al.
+2020; Triastcyn and Faltings 2019) or secure multi-party computation (MPC) (Truex et al.
+2019; Byrd and Polychroniadou 2020; Li et al.
+2020), to provide formal privacy guarantees.
 We now extend FedLog to be differentially private.
-Theorem 3.1.
-If the absolute value of features are clipped to $b$, FedLog messages satisfy $(\epsilon,\delta)$-DP with additive Gaussian noise ${\bf T}^{\prime}(\boldsymbol{\Phi}_{c},{\bf y}_{c}):={\bf T}(\boldsymbol{\Phi}_{c},{\bf y}_{c})+\mathcal{N}({\bf 0},\sigma^{2}{\bf I})$, where $\sigma=\sqrt{2(1+(m-1)*b^{2})\ln(1.25/\delta)}/{\epsilon}$.
+($\epsilon,\delta$)-DP protects clients’ privacy by adding noise to the shared information so that the adversaries cannot effectively tell if any record is included in the dataset (controlled by $\epsilon>0$) at most times (controlled by $10\leq\delta<$ 1) (Kerkouche et al.
+2021).
+Definition 3.
+A mechanism $M_{DP}$ satisfies $(\epsilon$, $\delta)$-DP if for any two datasets $\mathcal{D},\mathcal{D}^{\prime}$ that differ by only one record (i.e.
+$|(\mathcal{D}-\mathcal{D}^{\prime})\cup(\mathcal{D}^{\prime}-\mathcal{D})|=1$), and for any possible output $O\in Range(M_{DP})$,
+$\Pr_{O\sim M_{DP}(\mathcal{D})}\left[\log(\frac{\Pr[M_{DP}(\mathcal{D})=O]}{\Pr[M_{DP}(\mathcal{D}^{\prime})=O]})>\epsilon\right]<\delta$
+Intuitively, $(\epsilon,\delta)$-DP guarantees that the inner log ratio, considered as the information loss leaked to the adversaries, is bounded by the privacy budget $\epsilon$ with probability $\delta$.
+We add Gaussian noise to shared sufficient statistics as follows.
+Theorem 4.
+If the absolute value of features are clipped to $b$ and there are in total $k$ global update rounds, FedLog messages satisfy $(\epsilon,\delta)$-DP with additive Gaussian noise ${\bf T}^{\prime}(\boldsymbol{\Phi}_{c},{\bf y}_{c}):={\bf T}(\boldsymbol{\Phi}_{c},{\bf y}_{c})+\mathcal{N}({\bf 0},\sigma^{2}{\bf I})$, where $\sigma=\sqrt{8k(1+(m-1)*b^{2})\ln(e+\epsilon/\delta)}/{\epsilon}$.
 Proof.
-$L_{2}({\bf T})=\max_{\mathcal{D},\mathcal{D}^{\prime}}||{\bf T}(\mathcal{D})-{\bf T}(\mathcal{D}^{\prime})||_{2}=\max_{\boldsymbol{\phi},y}||{\bf T}(\boldsymbol{\phi},y)||_{2}=\sqrt{1+(m-1)*b^{2}}$
-See more details in Appendix A. ∎
+$\max_{\mathcal{D},\mathcal{D}^{\prime}}||{\bf T}(\mathcal{D})-{\bf T}(\mathcal{D}^{\prime})||_{2}$ $=\max_{\boldsymbol{\phi},y}||{\bf T}(\boldsymbol{\phi},y)||_{2}$
+$=\sqrt{1+(m-1)*b^{2}}$
+See more details in Appendix C. ∎
 When $n$ is large enough, the amount of noise (independent of $n$) becomes negligible to the model.
-Optionally, with secure MPC, it is sufficient to add such Gaussian noise once globally to the summation of all the messages each global update round.
+Optionally, with secure MPC or central DP, it is sufficient to add such Gaussian noise once globally to $\boldsymbol{\Phi}$ each global update round.
 The server then computes the posterior normally with the noisy sufficient statistics.
-4 Experiments
-4.1 Synthetic
+Table 1: Testing accuracy and communication cost reported for MNIST, CIFAR10, and CIFAR100.
+Accuracy reports the mean $\pm$ standard error of the testing accuracy over $10$ seeds.
+Higher is better.
+Communication cost reports the total message size transmitted between clients and the server.
+Lower is better.
+$\Uparrow$ denotes significantly higher results with $p<0.01$; $\Downarrow$ denotes significantly lower results with $p<0.01$.
+MNIST CIFAR10 CIFAR100
+accuracy comm cost accuracy comm cost accuracy comm cost
+FedAvg 89.76$\pm$0.69 3.45$\pm$0.02Gb 26.29$\pm$0.44 37.9$\pm$1.31Gb 13.34$\pm$0.15 69.1$\pm$0.47Gb
+LG-FedAvg 1 97.85$\pm$0.05 4.81$\pm$0.31Mb 86.57$\pm$0.29 0.26$\pm$0.02Gb 55.00$\pm$0.26 4.34$\pm$0.18Gb
+LG-FedAvg 2 98.18$\pm$0.06 0.16$\pm$0.07Gb 85.56$\pm$0.32 3.38$\pm$0.32Gb 54.90$\pm$0.24 9.53$\pm$0.53Gb
+FedPer 96.16$\pm$0.19 0.65$\pm$0.05Gb 83.54$\pm$0.40 27.7$\pm$1.50Gb 52.82$\pm$0.21 42.7$\pm$1.47Gb
+FedRep 95.51$\pm$0.29 36.3$\pm$3.38Mb 82.96$\pm$0.35 15.3$\pm$1.33Gb 48.70$\pm$0.29 11.0$\pm$0.40Gb
+CS-FL 79.65$\pm$1.22 0.35$\pm$0.01Gb 23.60$\pm$1.08 2.72$\pm$0.41Gb 4.52$\pm$0.15 13.7$\pm$0.22Gb
+FedBabu 86.30$\pm$1.04 2.25$\pm$0.02Gb 25.37$\pm$0.44 34.7$\pm$2.29Gb 9.70$\pm$0.16 59.3$\pm$0.31Gb
+FedProto 98.19$\pm$0.06 3.02$\pm$0.17Mb 87.37$\pm$0.26 0.18$\pm$0.01Gb 55.32$\pm$0.19 3.01$\pm$0.12Gb
+FedDBE 96.79$\pm$0.34 1.71$\pm$0.39Gb 72.77$\pm$0.79 38.3$\pm$0.95Gb 36.67$\pm$0.85 55.1$\pm$2.80Gb
+FedLog (ours) 98.15$\pm$0.05 3.18$\pm$0.31Mb 87.08$\pm$0.22 0.14$\pm$0.01Gb 56.46$\pm$0.27 2.38$\pm$0.09Gb
+FedLog-C (ours) 98.41$\pm$0.07 3.18$\pm$0.15Mb 87.57$\pm$0.25 0.11$\pm$0.01Gb 56.78$\pm$0.26 2.74$\pm$0.12Gb
+Experiments
+Synthetic
 We designed the following synthetic experiment to justify our claim that FedLog can learn universal local representation spaces without sharing local feature extractors, and argue why FedLog can converge faster than prior arts.
@@
 The results show: i) FedLog makes more progress in one global update round than other averaging based prior arts; ii) FedLog is resistant to over-fitting, as the difference between training and testing accuracy is small; iii) FedLog can learn universal local representations by fixing the last layer, even with different initialization and architectures of the feature extractor.
-Figure 1: Synthetic experiments.
-Dots are data points or local representations: dark green: client 0 class 0; light green: client 1 class 0; Dark blue: client 0 class 1; light blue: client 1 class 1.
-Dashed lines are linear separators.
-Accuracy results are averaged over 6 seeds.
-Table 1: Testing accuracy and communication cost reported for MNIST, CIFAR10, and CIFAR100.
-Accuracy reports the mean $\pm$ standard error of the testing accuracy over $10$ seeds.
-Higher is better.
-Communication cost reports the number of bits transmitted between one client and the server each global update round.
-Lower is better.
-$\Uparrow$ denotes significantly better results with $p<0.01$; $\Downarrow$ denotes significantly worse results with $p<0.01$.
-MNIST CIFAR10 CIFAR100
-accuracy comm cost accuracy comm cost accuracy comm cost
-FedAvg 89.76$\pm$0.69 698880 26.29$\pm$0.44 4102528 13.34$\pm$0.15 4684288
-LG-FedAvg 1 97.85$\pm$0.05 16320 86.57$\pm$0.29 64640 55.00$\pm$0.26 646400
-LG-FedAvg 2 98.18$\pm$0.06 529920 85.56$\pm$0.32 839040 54.90$\pm$0.24 1420800
-FedPer 96.16$\pm$0.19 168960 83.54$\pm$0.40 3263488 52.82$\pm$0.21 3263488
-FedRep 95.51$\pm$0.29 168960 82.96$\pm$0.35 3263488 48.70$\pm$0.29 3263488
-CS-FL 79.65$\pm$1.22 72088 23.60$\pm$1.08 423092 4.52$\pm$0.15 966172
-FedBabu 86.30$\pm$1.04 682560 25.37$\pm$0.44 4037888 9.70$\pm$0.16 4037888
-FedProto 98.19$\pm$0.06 16320 87.37$\pm$0.26 64640 55.32$\pm$0.19 646400
-FedDBE 96.79$\pm$0.34 700480 72.77$\pm$0.79 4108928 36.67$\pm$0.85 4690688
-FedLog (ours) 98.15$\pm$0.05 16320 87.08$\pm$0.22 64640 56.46$\pm$0.27 646400
-Table 2: Lower is better. $\Uparrow$ denotes significantly worse results with $p<0.01$.
-MNIST CIFAR10 CIFAR100
-rounds needed rounds needed rounds needed
-LG-FedAvg 1 5.9$\pm$0.38 40.3$\pm$3.09 67.1$\pm$2.76
-FedProto 3.7$\pm$0.21 28.5$\pm$2.05 46.6$\pm$1.90
-FedLog(ours) 3.9$\pm$0.38 21.5$\pm$2.23 36.8$\pm$1.39
-4.2 Communication Cost
-We conduct experiments on MNIST, CIFAR10, and CIFAR100, to show FedLog achieves better accuracy with less communication.
-We compare FedLog with the following baselines: i) FedAvg [2], which averages the whole model each global update round; ii) LG-FedAvg [3], which localizes bodies and averages heads.
-Two variants are reported: LG-FedAvg 1 which maintains one global layer and LG-FedAvg 2 which maintains two global layers; iii) FedPer [4], which localizes heads and averages bodies; iv) FedRep [5], which also localizes heads and averages bodies, but trains local heads and bodies separately; v) CS-FL [24], a model compression technique that compresses messages with the compressed sensing framework; vi) FedBabu [35], which averages bodies and never updates heads; vii) FedProto [12], which averages local feature representations by class and forces local models to learn similar representations; viii) FedDBE [36], which averages both bodies and heads but accelerates convergence by learning a domain representation bias.
+Communication Cost
+To show FedLog achieves better accuracy with less communication with non-i.i.d.
+clients, we conduct experiments on MNIST, CIFAR10, and CIFAR100.
+We compare FedLog and FedLog-C with the following baselines: i) FedAvg (McMahan et al.
+2017), which averages the whole model each global update round; ii) LG-FedAvg (Liang et al.
+2020), which localizes bodies and averages heads.
+Two variants are reported: LG-FedAvg 1 which maintains one global layer and LG-FedAvg 2 which maintains two global layers; iii) FedPer (Arivazhagan et al.
+2019), which localizes heads and averages bodies; iv) FedRep (Collins et al.
+2021), which also localizes heads and averages bodies, but trains local heads and bodies separately; v) CS-FL (Li, Li, and Varshney 2021), a model compression technique that compresses messages with the compressed sensing framework; vi) FedBabu (Oh, Kim, and Yun 2021), which averages bodies and never updates heads; vii) FedProto (Tan et al.
+2022b), which averages local feature representations by class and forces local models to learn similar representations; viii) FedDBE (Zhang et al.
+2024), which averages both bodies and heads but accelerates convergence by learning a domain representation bias.
 Convolutional neural networks of the same architecture and initialization is used for all the algorithms.
-FedPer and FedRep localizes the last two fully connected layers for less communication cost.
-We first distribute the training set evenly into $50$ (for MNIST) or $100$ (for CIFAR10 and CIFAR100) clients.
-Each client takes only $2$ classes to maximize the heterogeneity.
+FedPer and FedRep localizes the last two layers.
+We first distribute the training set into $(50,100,100)$ heterogeneous clients for (MNIST, CIFAR10, CIFAR100) respectively.
+Each client takes only $(2,2,10)$ classes.
 The testing set is distributed similarly to the clients, following the same distribution as the training data.
-Specially, for MNIST only, to simulate the difficult situation where clients do not have sufficient local data, we train on only $5\%$ of the training set, while test on the whole testing set.
-All the clients start from the same initialization, and all the algorithms are run for $100$ (for MNIST and CIFAR10) or $150$ (for CIFAR100) global update rounds.
+Specially, for MNIST only, to simulate the difficult situation where clients do not have sufficient local data, we train on $5\%$ of the training set, while test on the whole testing set.
+All the clients start from the same initialization, and all the algorithms are run for $(100,100,150)$ global update rounds.
 The testing accuracy is recorded after each round.
 Hyperparameters are optimized beforehand through grid search.
-See Appendix B for the model architectures, hyperparameters used by each algorithm and other experiment details.
-We report the mean and the standard error of the testing accuracy resulting from $10$ seeds in Table 1.
-We measured the statistical significance of the results compared to FedLog with one-tailed Wilcoxon signed-rank tests [37].
-We also report the communication cost of each global update round, namely the number of bits transmitted between one client and the server.
-The results show FedLog can achieve statistically significant accuracy improvement or at least comparable utility compared to prior arts, with the least communication cost as small as $2.3\%$, $1.6\%$, and $13.8\%$ of the communication cost of FedAvg for MNIST, CIFAR10, and CIFAR100 respectively.
-We further report how many global rounds are needed to reach a reasonable accuracy threshold in Table 2.
-The thresholds are 97%, 83%, and 53% respectively.
-These numbers are the largest integer of accuracy which all three algorithms have reached in 10 seeds.
-The results show FedLog needs less rounds to converge than LG-FedAvg and FedProto, thus saving more communication cost despite the cost per round is the same.
-4.3 Celeba
-We compare FedLog to LG-FedAvg and FedProto on large dataset Celeba [38] preprocessed by LEAF [39].
+See Appendix D for the model architectures, hyperparameters used by each algorithm and other experiment details.
+We report mean $\pm$ standard error of the testing accuracy resulting from $10$ seeds in Table 1.
+We measured the statistical significance of the results compared to FedLog-C with one-tailed Wilcoxon signed-rank tests (Wilcoxon 1992).
+Similarly, we report the total communication cost, namely the total size of messages transmitted between clients and the server, to reach a reasonable accuracy threshold (97%, 83%, and 53% respectively).
+If it is never reached, we stop counting at the round with the highest accuracy.
+These thresholds are the largest integer of accuracy which all competitive algorithms (FedLog, LG-FedAvg, FedProto) have reached in 10 seeds.
+The results show FedLog can achieve statistically significant accuracy improvement compared to prior arts, with the least communication cost as small as $0.09\%$, $0.29\%$, and $3.44\%$ of the communication cost of FedAvg for MNIST, CIFAR10, and CIFAR100 respectively.
+The closest baseline is FedProto, which has a slightly lower performance and requires more communication on CIFAR10 and CIFAR100 because it uses FedAvg instead of Bayesian inference for aggregation.
+Figure 2: From top to bottom: Celeba, flexible architecture, and differential privacy results. Colored area shows mean $\pm$ standard deviation of the accuracy.
+Celeba
+We compare FedLog and FedLog-C to LG-FedAvg and FedProto on the large dataset Celeba (Liu et al.
+2015) preprocessed by LEAF (Caldas et al.
+2018).
 We sampled 70838 images from 2360 clients.
 Each client represents a celebrity, and the local data are images of the same person.
-Sample inputs are shown in the left half of Fig.
+The task is to classify if the celebrity is smiling.
+MobileNetV2 (Sandler et al.
+2018) is implemented as the classifier model for all algorithms.
+To test how these algorithms perform under the situation where only a few communication rounds are available, we run every algorithm to optimum locally before global aggregation.
+The results are shown in the top graph of Fig.
 2.
-The task is to classify if the celebrity is smiling.
-MobileNetV2 [40] is implemented as the classifier model for all 3 algorithms.
-To test how these algorithms perform under the situation where only a few communication rounds are available, we run every algorithm to optimum locally before global aggregation.
-The results are shown in the right half of Fig.
-2.
-FedLog performs statistically significantly better than FedProto (71.19$\pm$0.45 v.s.
-69.98$\pm$0.16, $p=0.03$) and LG-FedAvg (66.84$\pm$0.19, $p=0.001$).
-4.4 Flexible Architecture
+FedLog-C performs statistically significantly better than FedProto (71.96$\pm$0.18 v.s.
+69.98$\pm$0.16, $p=0.001$), LG-FedAvg (66.84$\pm$0.19, $p=0.001$), and FedLog (71.19$\pm$0.45, $p=0.05$).
+Flexible Architecture
 We simulate the situation where clients have different computational resources on CIFAR100.
@@
 Different clients start from different local body initialization, but the global head is unified.
-Other experiment details are similar to Sec.
-4.2.
-We compare FedLog to LG-FedAvg 1 and FedProto.
-As shown in the left graph of Fig.
-3, FedLog converges faster than LG-FedAvg and FedProto.
-The accuracy of FedLog is also statistically significantly higher than the accuracy of FedProto (55.31$\pm$0.15 v.s.
-54.74$\pm$0.13, $p=0.002$) and LG-FedAvg (54.03$\pm$0.08, $p=0.001$).
-Figure 2: Left: sample Celeba input. Right: model accuracy. Colored area shows mean $\pm$ standard deviation of the accuracy.
-Figure 3: Left: flexible architecture results. Right: differential privacy trade-off. Colored area shows mean $\pm$ standard deviation of the accuracy.
-4.5 Differential Privacy
+We compare FedLog and FedLog-C to LG-FedAvg 1 and FedProto.
+As shown in the middle graph of Fig.
+2, FedLog-C and FedLog converges faster than LG-FedAvg and FedProto.
+The accuracy of FedLog-C is also statistically significantly higher than FedProto (55.86$\pm$0.11 v.s.
+54.74$\pm$0.13, $p=0.001$), LG-FedAvg (54.03$\pm$0.08, $p=0.001$), and FedLog (55.31$\pm$0.15, $p=0.001$).
+Differential Privacy
 We conduct experiments to show the trade-off between privacy budget $\epsilon$ and the accuracy of FedLog on CIFAR10.
-Experiment details are similar to Sec.
-4.2, except we add an activation function to clip the extracted features to $b=2$, and $\delta=0.01$.
-As shown in the right graph of Fig.
-3, the accuracy of differentially private FedLog quickly grows back to optimum when $\epsilon\geq 0.1$.
-This is a very strong privacy budget, and the results show FedLog performs well without sacrificing clients’ privacy.
-5 Conclusion
+We add an activation function to clip the extracted features to $b=2$.
+Following a common practice in FL (Wei et al.
+2020), we set $\delta=0.01$.
+As shown in the bottom graph of Fig.
+2, the accuracy of differentially private FedLog quickly grows back to optimum when $\epsilon\geq 0.5$, if the server is trusted or MPC is implemented so central DP is applicable.
+This is a strong privacy budget that shows FedLog performs well without sacrificing clients’ privacy.
+Otherwise, local DP can have more impact on the model utility, but the accuracy is still acceptable when $\epsilon\geq 5.0$.
+Limitations
+One limitation of FedLog is that the algorithm works only for classification.
+The model is solely designed to mimic cross-entropy loss and therefore a different loss function with a different model would be needed for regression.
+FedLog also assumes an exponential family distribution with a prior of the form $exp(-A(\boldsymbol{\eta}))$ and local transformed features distributed according to a mixture of Gaussians.
+However the mixture of Gaussian assumption is mitigated in FedLog-C by introducing an auxiliary loss that helps satisfy the assumption.
+A benefit of the exponential family and mixture of Gaussian assumptions is that the data summaries are provable sufficient statistics (Thm.
+2) and we obtain a closed form solution for the normalization constant (Eq.
+11).
+Conclusion
 We proposed FedLog that shares local data summaries instead of model parameters.
-FedLog assumes an exponential family model on local data, and learns a global linear separator with the summation of sufficient statistics.
+FedLog assumes an exponential family model on local representations, and learns a global linear separator with the summation of sufficient statistics.
 FedLog can learn universal local representations without sharing the bodies.
@@
 It is also effective with flexible architectures and formal DP guarantees.
-One limitation of FedLog is the assumption that the embedding of the data (just prior to the classification head) follows a mixture of Gaussians.
-Relaxing this assumption in future work could further improve the performance of the approach.
-Acknowledgments and Disclosure of Funding
+For future work, it would be interesting to generalize FedLog to regression.
+Ackowledgment
 Resources used in preparing this research at the University of Waterloo were provided by Huawei Canada, the province of Ontario and the government of Canada through CIFAR and companies sponsoring the Vector Institute.
-Appendix A Differential Privacy
-It has been shown that FL algorithms that share large model parameters do not prevent privacy attacks through weight manipulation, GAN-based reconstruction, and large model memorization effects [41, 42].In our case, We argue that FedLog, which shares summations of sufficient statistics only, avoid these pitfalls closely related to model parameter sharing.
+Appendix A A. Other Related Works
+Communication efficient FL.
+Being orthogonal to FRL, communication efficient FL aims to directly decrease the communication overhead with optimization algorithms, client selection, and model compression (Wen et al.
+2023).
+First, since local training epochs affect the rounds of global communication needed (McMahan et al.
+2017), researchers proposed different local optimization methods to reduce the communication rounds (Liu et al.
+2020; Wu and Wang 2021; Wu et al.
+2022).
+Second, some clients may contribute more to the global model, or are faster when uploading parameters.
+Thus, the global learning process can be accelerated by carefully selecting clients that meet these criteria (Liu et al.
+2021b; Deng et al.
+2021; Lai et al.
+2021; Du, Xiao, and Guo 2022).
+Additionally, the size of transmitted messages can be directly decreased by reducing or compressing the model parameters (Lu et al.
+2020; Li and Xiao 2021; Li, Li, and Varshney 2021).
+However, most such algorithms are efficient at the cost of model accuracy (Cai et al.
+2022).
+Bayesian FL.
+Other Bayesian models have been explored to represent distributions over models and predictions in FL.
+The challenge is in the aggregation of the local posteriors.
+Various techniques have been proposed including personalized GPs (Achituve et al.
+2021), posterior averaging (Al-Shedivat et al.
+2020), online Laplace approximation (Liu et al.
+2021a), Thompson sampling (Dai, Low, and Jaillet 2020), MCMC sampling (Vono et al.
+2022).
+These Bayesian FL techniques tend to emphasize calibration, approximating the posterior, or even different tasks.
+There is not much in common between them and our approach despite the use of Bayes theorem.
+Appendix B B. Sufficiency
+This section discusses formal definitions of sufficient and other statistics, based on Chapter 6 of (Casella and Berger 2015).
+Sufficient Statistics.
+Definition A.5 (6.2.1 in (Casella and Berger 2015)).
+A statistic ${\bf T}({\bf x})$ is a sufficient statistic for $\boldsymbol{\eta}$ if the conditional distribution of the sample ${\bf x}$ given the value of ${\bf T}({\bf x})$ does not depend on $\boldsymbol{\eta}$.
+Intuitively, a sufficient statistic captures all the information about $\boldsymbol{\eta}$ in ${\bf X}$. In the case of the exponential family, the following theorem applies:
+Theorem A.6 (6.2.10 in (Casella and Berger 2015)).
+Let ${\bf x}_{1},\cdots,{\bf x}_{n}$ be i.i.d.
+observations from an exponential family whose p.d.f.
+is given by $\Pr({\bf x}|\boldsymbol{\eta})=h({\bf x})\exp(\boldsymbol{\eta}^{\top}{\bf T}({\bf x})-A(\boldsymbol{\eta}))$, then ${\bf T}({\bf x}_{1},\cdots,{\bf x}_{n})=\sum_{i=1}^{n}{\bf T}({\bf x}_{i})$ is a sufficient statistic for $\boldsymbol{\eta}$.
+Minimal Sufficient Statistics.
+Definition A.7 (6.2.11 in (Casella and Berger 2015)).
+A sufficient statistic $\bf T(x)$ is minimal if, for any other sufficient statistic $\bf T^{\prime}(x)$, $\exists$ a function $h$ such that $h({\bf T^{\prime}(x)})={\bf T(x)}$.
+A minimal sufficient statistic achieves the greatest possible data reduction for a sufficient statistic. Whether a sufficient statistic is minimal can be verified by the following theorem.
+Theorem A.8 (6.2.13 in (Casella and Berger 2015)).
+If for all sample points $\bf x,x^{\prime}$ from the distribution with p.d.f.
+$\Pr({\bf x}|\boldsymbol{\eta})$, $\frac{\Pr({\bf x}|\boldsymbol{\eta})}{\Pr({\bf x^{\prime}}|\boldsymbol{\eta})}$ is independent of $\boldsymbol{\eta}$ iff $\bf T(x)=T(x^{\prime})$, then $\bf T(x)$ is minimal sufficient.
+Ancillary Statistics.
+Definition A.9 (6.2.16 in (Casella and Berger 2015)).
+A statistic $\bf S(x)$ is an ancillary statistic if it is independent of the parameters $\boldsymbol{\eta}$.
+As shown by the definition, an ancillary statistic on its own contains no information about the parameters $\boldsymbol{\eta}$.
+Complete Statistics.
+Definition A.10 (6.2.21 in (Casella and Berger 2015)).
+A family of distributions is called complete if $\mathbb{E}_{\boldsymbol{\eta}}g({\bf T})=0,\forall\boldsymbol{\eta}$ implies $\Pr_{\boldsymbol{\eta}}(g({\bf T})=0)=1,\forall\boldsymbol{\eta}$.
+Then $\bf T(x)$ is a complete statistic.
+This definition is less intuitive and harder to interpret. We skip the details, but focus on the following theorems.
+Theorem A.11 (6.2.25 in (Casella and Berger 2015)).
+Let ${\bf x}_{1},\cdots,{\bf x}_{n}$ be i.i.d.
+observations from an exponential family whose p.d.f.
+is given by $\Pr({\bf x}|\boldsymbol{\eta})=h({\bf x})\exp(\boldsymbol{\eta}^{\top}{\bf T}({\bf x})-A(\boldsymbol{\eta}))$, then ${\bf T}({\bf x}_{1},\cdots,{\bf x}_{n})=\sum_{i=1}^{n}{\bf T}({\bf x}_{i})$ is a complete statistic for $\boldsymbol{\eta}$ if support of $\boldsymbol{\eta}$ is open.
+Theorem A.12 (Basu’s Theorem, 6.2.24 in (Casella and Berger 2015)).
+If $\bf T(x)$ is a complete and minimal sufficient statistic, then $\bf T(x)$ is independent of every ancillary statistic.
+Proof of Our Claim
+We now prove our claim made in the main paper. We copy Theorem 2 from the main paper here.
+Theorem 2.
+If $({\boldsymbol{\phi}_{1}},y_{1}),({\boldsymbol{\phi}_{2}},y_{2}),\cdots,({\boldsymbol{\phi}_{n}},y_{n})$ are i.i.d.
+samples from the exponential family with ${\bf T}({\boldsymbol{\phi}},y)=\boldsymbol{\phi}\otimes{\bf e}_{y}$, then ${\bf T}(({\boldsymbol{\phi}_{1}},y_{1}),\cdots,({\boldsymbol{\phi}_{n}},y_{n}))=\sum_{i=1}^{n}{\boldsymbol{\phi}_{i}}\otimes{\bf e}_{y_{i}}$ is a minimal sufficient statistic independent of every ancillary statistic.
+Proof.
+  1. 1.
+    ${\bf T}(({\boldsymbol{\phi}_{1}},y_{1}),({\boldsymbol{\phi}_{2}},y_{2}),\cdots,({\boldsymbol{\phi}_{n}},y_{n}))$ is sufficient by Thm. A.6.
+  2. 2.
+    ${\bf T}(({\boldsymbol{\phi}_{1}},y_{1}),({\boldsymbol{\phi}_{2}},y_{2}),\cdots,({\boldsymbol{\phi}_{n}},y_{n}))$ is complete by Thm.
+A.11, since $\boldsymbol{\eta}\in\mathbb{R}^{m*n_{class}}$ is open.
+  3. 3.
+    Let
+    $R$ $=\frac{\Pr(({\boldsymbol{\phi}_{1}},y_{1}),({\boldsymbol{\phi}_{2}},y_{2}),\cdots,({\boldsymbol{\phi}_{n}},y_{n})|\boldsymbol{\eta})}{\Pr(({\boldsymbol{\phi}^{\prime}_{1}},y^{\prime}_{1}),({\boldsymbol{\phi}^{\prime}_{2}},y’_{2}),\cdots,({\boldsymbol{\phi}^{\prime}_{n}},y^{\prime}_{n})|\boldsymbol{\eta})}$
+    $=\exp(\boldsymbol{\eta}^{\top}(\sum_{i=1}^{n}{\boldsymbol{\phi}_{i}}\otimes{\bf e}_{y_{i}}-\sum_{i=1}^{n}{\boldsymbol{\phi}^{\prime}_{i}}\otimes{\bf e}_{y^{\prime}_{i}}))$
+    ${\bf T}(({\boldsymbol{\phi}_{1}},y_{1}),({\boldsymbol{\phi}_{2}},y_{2}),\cdots,({\boldsymbol{\phi}_{n}},y_{n}))$ is also minimal by Thm.
+A.8, since $R$ is independent of $\boldsymbol{\eta}$ if and only if
+    ${\bf T}(({\boldsymbol{\phi}_{1}},y_{1}),\cdots,({\boldsymbol{\phi}_{n}},y_{n}))={\bf T}(({\boldsymbol{\phi}^{\prime}_{1}},y^{\prime}_{1}),\cdots({\boldsymbol{\phi}^{\prime}_{n}},y^{\prime}_{n}))$
+  4. 4.
+    ${\bf T}(({\boldsymbol{\phi}_{1}},y_{1}),({\boldsymbol{\phi}_{2}},y_{2}),\cdots,({\boldsymbol{\phi}_{n}},y_{n}))$ is independent of every ancillary statistic by Thm. A.12.
+Appendix C C. Differential Privacy
+It has been shown that FL algorithms that share large model parameters do not prevent privacy attacks through weight manipulation, GAN-based reconstruction, and large model memorization effects (Boenisch et al.
+2021; Mothukuri et al.
+2021).
+We argue that FedLog, which shares summations of sufficient statistics only, avoid these pitfalls closely related to model parameter sharing.
 Since “addition” is a non-invertible function, malicious attackers cannot recover features of individual data points.
@@
 To further guarantee users’ privacy formally, we now extend FedLog to be differentially private.
-($\epsilon,\delta$)-DP protects clients’ privacy by adding noise to the shared information so that the adversaries cannot effectively tell if any record is included in the dataset (controlled by $\epsilon>0$) at most times (controlled by $10\leq\delta<$ 1) [43].
+($\epsilon,\delta$)-DP protects clients’ privacy by adding noise to the shared information so that the adversaries cannot effectively tell if any record is included in the dataset (controlled by $\epsilon>0$) at most times (controlled by $10\leq\delta<$ 1) (Kerkouche et al.
+2021).
 Definition A.9.
 A mechanism $M_{DP}$ satisfies $(\epsilon$, $\delta)$-DP if for any two datasets $\mathcal{D},\mathcal{D}^{\prime}$ that differ by only one record (i.e.
-$|\mathcal{D}-\mathcal{D}^{\prime}\cup\mathcal{D}^{\prime}-\mathcal{D}|=1$), and for any possible output $O\in Range(M_{DP})$,
+$|(\mathcal{D}-\mathcal{D}^{\prime})\cup(\mathcal{D}^{\prime}-\mathcal{D})|=1$), and for any possible output $O\in Range(M_{DP})$,
 $\Pr_{O\sim M_{DP}(\mathcal{D})}\left[\log(\frac{\Pr[M_{DP}(\mathcal{D})=O]}{\Pr[M_{DP}(\mathcal{D}^{\prime})=O]})>\epsilon\right]<\delta$
@@
 A commonly used mechanism is to add Gaussian noise to $f(\mathcal{D})$:
-Theorem A.3 (e.g. [43]).
-A Gaussian Mechanism $\mathcal{GM}$ is $(\epsilon,\delta)$-DP with information revealing function $f$: $\mathcal{GM}(\mathcal{D})=f(\mathcal{D})+\mathcal{N}({\bf 0},\sigma^{2}{\bf I})$, where $\sigma={L_{2}(f)}\sqrt{2\ln(1.25/\delta)}/{\epsilon}$.
+Theorem A.11 ((Kairouz, Oh, and Viswanath 2015)).
+For real-valued queries with sensitivity $L_{2}(f)>0$, the mechanism that adds Gaussian noise with standard deviation $\sqrt{8k\ln(e+\epsilon/\delta)}L_{2}(f)/{\epsilon}$ satisfies $(\epsilon,\delta)$-differential privacy under $k$-fold adaptive composition, $\forall\epsilon>0,\delta\in(0,1]$.
 In FedLog, the only private information shared by clients is the summation of statistics ${\bf T}(\boldsymbol{\Phi}_{c},{\bf y}_{c})$, a vector of size $n_{class}*m$.
@@
 $g(x):=\begin{cases}b,&\text{if }x>b\\ -b,&\text{if }x<-b\\ x,&\text{otherwise}\end{cases}$ (16)
-Then we can calculate the $L_{2}$ sensitivity as follows
-$\begin{split}L_{2}({\bf T})&=\max_{\mathcal{D},\mathcal{D}^{\prime}}||{\bf T}(\mathcal{D})-{\bf T}(\mathcal{D}^{\prime})||_{2}\\ &=\max_{\boldsymbol{\phi},y}||{\bf T}(\boldsymbol{\phi},y)||_{2}\\ &=\sqrt{1+(m-1)*b^{2}}\end{split}$ (17)
-Appendix B Experiment Details
+We now prove Theorem 4 from the main paper.
+Theorem 4.
+If the absolute value of features are clipped to $b$ and there are in total $k$ global update rounds, FedLog messages satisfy $(\epsilon,\delta)$-DP with additive Gaussian noise ${\bf T}^{\prime}(\boldsymbol{\Phi}_{c},{\bf y}_{c}):={\bf T}(\boldsymbol{\Phi}_{c},{\bf y}_{c})+\mathcal{N}({\bf 0},\sigma^{2}{\bf I})$, where $\sigma=\sqrt{8k(1+(m-1)*b^{2})\ln(e+\epsilon/\delta)}/{\epsilon}$.
+Proof.
+We calculate the $L_{2}$ sensitivity as follows
+$L_{2}({\bf T})$ $=\max_{\mathcal{D},\mathcal{D}^{\prime}}||{\bf T}(\mathcal{D})-{\bf T}(\mathcal{D}^{\prime})||_{2}$ (17)
+$=\max_{\boldsymbol{\phi},y}||{\bf T}(\boldsymbol{\phi},y)||_{2}$ (18)
+$=||[1,b,b,\cdots,b,0,0,\cdots,0]||_{2}$ (19)
+$=\sqrt{1+(m-1)*b^{2}}$ (20)
+Eq.
+17 equals to Eq.
+18 due to our definition of neighbouring datasets (adding or removing one record).
+Eq.
+18 equals to Eq.
+19 since one of our features is always 1, and there are at most $m-1$ other non-zero entries with maximum value of $b$.
+Finally, we apply Thm. A.11 to get $\sigma=\sqrt{8k(1+(m-1)*b^{2})\ln(e+\epsilon/\delta)}/{\epsilon}$. ∎
+Appendix D D. Experiment Details
 Communication cost, flexible architecture, and differential privacy experiments are run on 1 NVIDIA T4 GPU with 16GB RAM.
@@
 We make our code public in the supplementary materials, where further details can be found.
-Appendix C Licences
+Appendix E E. Licences
 Yann LeCun and Corinna Cortes hold the copyright of MNIST dataset, which is a derivative work from original NIST datasets.
 MNIST dataset is made available under the terms of the Creative Commons Attribution-Share Alike 3.0 license.
-The CIFAR-10 and CIFAR-100 are labeled subsets of the 80 million tiny images dataset. They were collected by Alex Krizhevsky, Vinod Nair, and Geoffrey Hinton, made public at the CIFAR homepage.
-The CelebA dataset is available for non-commercial research purposes only. See CelebA homepage for the full agreement.
+The CIFAR-10 and CIFAR-100 are labeled subsets of the 80 million tiny images dataset.
+They were collected by Alex Krizhevsky, Vinod Nair, and Geoffrey Hinton, made public at https://www.cs.toronto.edu/~kriz/cifar.html, the CIFAR homepage.
+The CelebA dataset is available for non-commercial research purposes only. See https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html, the CelebA homepage for the full agreement.
 MNIST CIFAR10 CIFAR100
@@
 local epochs 1
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: We achieved the smallest message size and fastest convergence without losing accuracy as shown in the experiment section. We illustrated the trade-off of the DP version.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: See section 3.2 and the conclusion.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: Full proof of theorem 3.1 is in Appendix A.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: Experiment details are described along with the results and in Appendix B.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: See supplemental material.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Full details are provided with the public code.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: One-tailed Wilcoxon signed-rank tests are done for all experiments.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: See Appendix B.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification:
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [No]
-  49.
-Justification: This paper presents work whose goal is to advance the field of Machine Learning.
-There could be potential societal consequences of our work, none which we feel must be specifically highlighted here.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: The paper poses no such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: Datasets used for experiments are properly cited. See Appendix C for the licenses.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: The paper does not release new assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.